### v1

# The torus plateau for the high-dimensional Ising model

Abstract
We consider the Ising model on a $d$-dimensional discrete torus of volume $r^{d}$, in dimensions $d>4$ and for large $r$, in the vicinity of the infinite-volume critical point $\beta_{c}$. We prove that for $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ (with a suitable constant) the susceptibility is bounded above and below by multiples of $r^{d/2}$, and that the two-point function has a “plateau” in the sense that it decays like $|x|^{-(d-2)}$ when $|x|$ is small relative to the volume but for larger $|x|$ it levels off to a constant value of order $r^{-d/2}$. We also prove that at $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ the renormalised coupling constant is nonzero, which implies a non-Gaussian limit for the average spin. The proof relies on near-critical estimates for the infinite-volume two-point function obtained recently by Duminil-Copin and Panis, and builds upon a strategy proposed by Papathanakos. The random current representation of the Ising model plays a central role in our analysis.
1 Introduction and main results
1.1 Introduction
The Ising model has been studied for a century as a fundamental example of a phase transition in statistical mechanics. We are interested here in the finite-size scaling of the Ising model on a finite box in the Euclidean lattice $\mathbb{Z}^{d}$ in dimensions $d>4$, particularly in the case of periodic boundary conditions which make the box into a torus.
For $d\geq 1$ and an integer $r\geq 3$, we write $\Lambda_{r}$ for the discrete box $[-r/2,r/2)^{d}\cap\mathbb{Z}^{d}$ of volume $r^{d}$, and write $\mathbb{T}_{r}=(\mathbb{Z}/r\mathbb{Z})^{d}$ for the $d$-dimensional discrete torus of period $r$. We study large $r$, and seek estimates that are uniform in large $r$. For convenience, we often identify a point $x\in\mathbb{T}_{r}$ with its representative in $\Lambda_{r}$.
Notation. We write $f\lesssim g$ to denote the existence of a constant $C>0$ such that $f\leq Cg$, write $f\asymp g$ to mean $f\lesssim g$ and $g\lesssim f$, and write $f\sim g$ to denote $\lim f/g=1$. To avoid division by zero, for $x\in\mathbb{R}^{d}$ we write $|\!|\!|x|\!|\!|=1\vee\|x\|_{\infty}$.
Let $\mathcal{G}=(V,E)$ be a finite graph with vertex set $V$ and edge set $E$. The Ising model on $\mathcal{G}$ is a family of probability measures on spin configurations $\sigma:V\to\{-1,+1\}$, defined using the Hamiltonian
$H^{\mathcal{G}}(\sigma)=-\sum_{xy\in E}\sigma_{x}\sigma_{y}.$ (1.1)
The expectation of a function $F$ of spin configurations, at inverse temperature $\beta>0$, is defined by
$\langle F\rangle_{\beta}^{\mathcal{G}}=\frac{1}{Z^{\mathcal{G}}_{\beta}}\sum_{\sigma}F(\sigma)e^{-\beta H^{\mathcal{G}}(\sigma)},$ (1.2)
where $Z^{\mathcal{G}}_{\beta}=\sum_{\sigma}e^{-\beta H^{\mathcal{G}}(\sigma)}$ is the partition function. Free boundary conditions (FBC) on the box correspond to $\mathcal{G}=\Lambda_{r}$ with edges connecting nearest neighbours in $\Lambda_{r}$. Periodic boundary conditions (PBC) correspond to $\mathcal{G}=\mathbb{T}_{r}$ with edges connecting nearest neighbours in $\mathbb{T}_{r}$, which include all edges of ${\Lambda_{r}}$ and additional edges that join opposite sides of $\Lambda_{r}$. The two-point functions are defined by
$\tau^{\Lambda_{r}}_{\beta}(x,y)=\langle\sigma_{x}\sigma_{y}\rangle_{\beta}^{\Lambda_{r}},\qquad\tau^{\mathbb{T}_{r}}_{\beta}(x,y)=\langle\sigma_{x}\sigma_{y}\rangle_{\beta}^{\mathbb{T}_{r}},$ (1.3)
and the finite-volume susceptibilities are defined by
$\chi^{\Lambda_{r}}(\beta)=\sum_{x\in\Lambda_{r}}\tau^{\Lambda_{r}}_{\beta}(0,x),\qquad\chi^{\mathbb{T}_{r}}(\beta)=\sum_{x\in\mathbb{T}_{r}}\tau^{\mathbb{T}_{r}}_{\beta}(0,x).$ (1.4)
It follows from the Griffiths inequalities that $\tau_{\beta}^{\Lambda_{r}}$ is nonnegative and increasing in $r$, so that the infinite-volume two-point function $\tau_{\beta}(x,y)=\lim_{r\to\infty}\tau_{\beta}^{\Lambda_{r}}(x,y)$ and infinite-volume susceptibility $\chi(\beta)=\lim_{r\to\infty}\chi^{\Lambda_{r}}(\beta)$ exist in $[0,\infty]$. Moreover, it is known that there is a critical inverse temperature $\beta_{c}$, below which $\tau_{\beta}(x,y)$ decays exponentially as $\lvert x-y\rvert\to\infty$ and ${\chi(\beta)}$ is finite. The susceptibility is infinite at $\beta_{c}$, and from [2] we know that for $d>4$ and $\beta<\beta_{c}$,
$\chi(\beta)\asymp\frac{1}{\beta_{c}-\beta}.$ (1.5)
An essential ingredient in our analysis is the very recent result from [17] that, for $d>4$, there are positive constants $c_{0},C_{0}$ such that for all $\beta\leq\beta_{c}$,
$\tau_{\beta}(0,x)$ $\leq\frac{C_{0}}{|\!|\!|x|\!|\!|^{d-2}}e^{-c_{0}(\beta_{c}-\beta)^{1/2}\lVert x\rVert_{\infty}}$ $\qquad\qquad(x\in\mathbb{Z}^{d}),$ (1.6)
$\tau_{\beta}(0,x)$ $\geq\frac{c_{0}}{|\!|\!|x|\!|\!|^{d-2}}$ $(\lVert x\rVert_{\infty}\leq c_{0}(\beta_{c}-\beta)^{-1/2}).$ (1.7)
For the spread-out Ising model in dimensions $d>4$, or for the nearest-neighbour model in sufficiently high dimensions, more precise asymptotics for the critical two-point function have been proved using the lace expansion [39].
The influence of boundary conditions on finite-size critical behaviour in dimensions $d\geq 4$ has been studied extensively in the physics literature for Ising and related models, e.g., [11, 14, 15, 31, 32, 41, 42]. In particular, it has been observed numerically and via physics scaling arguments that, for $d>4$,
$\tau^{\Lambda_{r}}_{\beta_{c}}(0,x)\asymp\frac{1}{|\!|\!|x|\!|\!|^{d-2}},\quad\tau^{\mathbb{T}_{r}}_{\beta_{c}}(0,x)\asymp\frac{1}{|\!|\!|x|\!|\!|^{d-2}}+\frac{1}{r^{d/2}},\qquad\chi^{\Lambda_{r}}(\beta_{c})\asymp r^{2},\quad\chi^{\mathbb{T}_{r}}(\beta_{c})\asymp r^{d/2}.$ (1.8)
(The claim for $\tau^{\Lambda_{r}}_{\beta_{c}}(0,x)$ is for $x$ not too close to the boundary of the box.) The constant term $r^{-d/2}$ in $\tau^{\mathbb{T}_{r}}_{\beta_{c}}(0,x)$ is referred to as the plateau, and it is responsible for the larger susceptibility for PBC compared to FBC.
In recent years, the plateau for PBC has been proven to exist for simple random walk in dimensions $d>2$ [40, 15], self-avoiding walk for $d>4$ [40, 30], and percolation for $d>6$ [27, 28]. A general theory of the effect of FBC vs PBC in the setting of the weakly-coupled hierarchical $n$-component $|\varphi|^{4}$ model in dimensions $d\geq 4$ was developed in [33, 37], and we comment further on this below.
For the Ising model in dimensions $d>4$, the absence of a plateau for FBC was proven in [13]. For PBC, incomplete steps were taken in [36] towards proving existence of the plateau. In this paper, we build on the method of [36] to prove (1.8) for PBC in dimensions $d>4$. We also prove the non-vanishing of the renormalised coupling constant under PBC, indicating a non-Gaussian limit for the average spin. This can be contrasted with the Gaussian limit under FBC obtained in [2, 20, 4, 35].
We emphasise that we do not use the lace expansion in any way, and our results do not require a small parameter. The input from the lace expansion used to establish plateau theorems for self-avoiding walk and percolation in [40, 30, 27, 28] is replaced here by the near-critical bounds (1.6)–(1.7).
1.2 Main results
Our first theorem indicates that for $\beta$ in an interval slightly below $\beta_{c}$, the torus and infinite-volume two-point functions are related by a plateau term $\chi(\beta)/r^{d}$.
Theorem 1.1.
Let $d>4$. There are constants $c_{i}>0$, depending on $d$ but not on $\beta$, such that
$\tau_{\beta}^{\mathbb{T}_{r}}(0,x)\leq\tau_{\beta}(0,x)+c_{1}\frac{\chi(\beta)}{r^{d}}$ (1.9)
for all $r\geq 3$, all $\beta\in(0,\beta_{c})$, and all $x\in{\mathbb{T}_{r}}$, and
$\tau_{\beta}^{\mathbb{T}_{r}}(0,x)\geq\tau_{\beta}(0,x)+c_{2}\frac{\chi(\beta)}{r^{d}}$ (1.10)
for all $r\geq 3$, all $\beta\in[\beta_{c}-c_{3}r^{-2},\beta_{c}-c_{4}r^{-d/2}]$ and all $\|x\|_{\infty}\geq c_{5}$. Moreover, the constant $c_{3}$ is given by $c_{3}=\frac{4}{9}c_{0}^{2}$ with $c_{0}$ the constant of (1.7).
Note that for the lower bound, the interval of $\beta$ values is nonempty when $r$ is sufficiently large, since $d>4$. In this case, we write
$\beta_{*}=\beta_{*}(r)=\beta_{c}-c_{4}r^{-d/2}.$ (1.11)
At $\beta_{*}$, we have $\chi(\beta_{*})\asymp r^{d/2}$ by (1.5), so (1.9)–(1.10) identify a plateau term in $\tau_{\beta_{*}}^{\mathbb{T}_{r}}$ of order $r^{-d/2}$. The monotonicity in $\beta$ then implies that the plateau term persists for larger $\beta$, in particular for $\beta_{c}$.
Corollary 1.2.
Let $d>4$ and $r$ be large. Then
$\tau_{\beta_{*}}^{\mathbb{T}_{r}}(0,x)\asymp|\!|\!|x|\!|\!|^{-(d-2)}+r^{-d/2}$ (1.12)
for all $x\in\mathbb{T}_{r}$. In particular, $\tau_{\beta}^{\mathbb{T}_{r}}(0,x)\gtrsim|\!|\!|x|\!|\!|^{-(d-2)}+r^{-d/2}$ for all $\beta\geq\beta_{*}$.
Summation of (1.12) over $x\in\mathbb{T}_{r}$ immediately gives the following bounds on the susceptibility, because $\sum_{x\in{\mathbb{T}_{r}}}|\!|\!|x|\!|\!|^{-(d-2)}=O(r^{2})=o(r^{d/2})$ in dimensions $d>4$.
Corollary 1.3.
For $d>4$ and $r$ large,
$\chi^{\mathbb{T}_{r}}(\beta_{*})\asymp r^{d/2}.$ (1.13)
In particular, $\chi^{\mathbb{T}_{r}}(\beta)\gtrsim r^{d/2}$ for all $\beta\geq\beta_{*}$.
Let $S_{r}=r^{-d}\sum_{x\in\mathbb{T}_{r}}\sigma_{x}$ denote the average spin on the torus. We define the renormalised coupling constant by
$g^{\mathbb{T}_{r}}(\beta)=-\frac{\langle S_{r}^{4}\rangle_{\beta}^{\mathbb{T}_{r}}-3(\langle S_{r}^{2}\rangle_{\beta}^{\mathbb{T}_{r}})^{2}}{(\langle S_{r}^{2}\rangle_{\beta}^{\mathbb{T}_{r}})^{2}}.$ (1.14)
By the Lebowitz inequality, the above numerator is non-positive, so $g^{\mathbb{T}_{r}}(\beta)\geq 0$. By the second Griffiths inequality, $\langle S_{r}^{4}\rangle_{\beta}^{\mathbb{T}_{r}}\geq(\langle S_{r}^{2}\rangle_{\beta}^{\mathbb{T}_{r}})^{2}$, so $g^{\mathbb{T}_{r}}(\beta)\leq 2$. The following theorem shows that $g^{\mathbb{T}_{r}}(\beta_{*})$ is bounded away from zero. This indicates a non-Gaussian limit for the average field at $\beta=\beta_{*}$: for a Gaussian random variable the numerator in (1.14) is zero. This is in contrast to the situation at (and below) $\beta_{c}$ with free boundary conditions [2, 4, 20, 35], where the limit is Gaussian. We do not have a proof that $g^{\mathbb{T}_{r}}$ is increasing in $\beta$, but see (1.20) below.
Theorem 1.4.
Let $d>4$. There is a constant $c_{g}>0$ such that for all $r$ large,
$0<c_{g}\leq g^{\mathbb{T}_{r}}(\beta_{*})\leq 2$ (1.15)
Our lower bounds on the susceptibility in Corollary 1.3, on the two-point function in Corollary 1.2, and on the renormalised coupling constant in Theorem 1.4, are claimed in [36], but the arguments put forth in [36] do not constitute a proof. Nevertheless, a useful strategy is proposed in [36], and we build upon that strategy in our work.
1.3 Conjectured asymptotic formulas
A precise and detailed account of the effect of boundary conditions on finite-size scaling is given in [33, 37], via a rigorous renormalisation group analysis. The results of [33, 37] are proved for the weakly-coupled $n$-component $|\varphi|^{4}$ model on the $d$-dimensional hierarchical lattice for $d\geq 4$ and for all $n\geq 1$, but, on the basis of universality, the results obtained there for the $1$-component model are conjectured to apply also to all $1$-component spin models in the universality class of the nearest-neighbour Ising model. An antecedent of the conjectures goes back to [12] (see also [43, Chapter 32]). The conjectures present predictions for a specific and universal non-Gaussian limit for the average field, and precise asymptotic amplitudes for the susceptibility, the plateau, and the renormalised coupling constant, throughout a critical window of width $r^{-d/2}$ (with a logarithmic correction for $d=4$) centred at $\beta_{c}$. We define the window scale and the large-field scale, respectively, by
$w_{r}=\begin{cases}a_{4}(\log r)^{-1/6}r^{-2}&(d=4)\\ a_{d}\,r^{-d/2}&(d>4),\end{cases}\qquad h_{r}=\begin{cases}b_{4}(\log r)^{1/4}r^{-1}&(d=4)\\ b_{d}\,r^{-d/4}&(d>4),\end{cases}$ (1.16)
with suitably chosen constants $a_{d},b_{d}>0$. We also define the universal profile $f:\mathbb{R}\to(0,\infty)$ by
$f(s)=\int_{\mathbb{R}}x^{2}\mathrm{d}\sigma_{s},\qquad\mathrm{d}\sigma_{s}=\frac{e^{-\frac{1}{4}x^{4}+\frac{s}{2}x^{2}}\mathrm{d}x}{\int_{\mathbb{R}}e^{-\frac{1}{4}x^{4}+\frac{s}{2}x^{2}}\mathrm{d}x}.$ (1.17)
The function $f$ is strictly monotone increasing, and satisfies $f(s)\sim|s|^{-1}$ as $s\to-\infty$ and $f(s)\sim s$ as $s\to\infty$ (see [33, Section 1.5.2] where $f_{1}(s)=f(-s)$). The following conjecture proposes a precise non-Gaussian limit for the average field $S_{r}$, and a precise role for the universal profile $f$ in the critical finite-size scaling of the susceptibility, the two-point function, and the renormalised coupling constant for the Ising model. The conjecture is a special case ($n=1$) of conjectures from [33, 37] for $n$-component models for all $n\geq 0$ ($n=0$ corresponds to the self-avoiding walk). The profile $f$ also arises in the Curie–Weiss model; see [18, Theorem 3.2].
Conjecture 1.5 (Non-gaussian limit and universal profile).
For $d\geq 4$, $s\in\mathbb{R}$, and $\beta=\beta_{c}+sw_{r}$, as $r\to\infty$ the rescaled average field converges in distribution to a real-valued random variable with the non-Gaussian distribution $\mathrm{d}\sigma_{s}$:
$h_{r}^{-1}S_{r}\Rightarrow\mathrm{d}\sigma_{s}\quad\text{under $\langle\cdot\rangle_{\beta_{c}+sw_{r}}^{{\mathbb{T}_{r}}}$.}$ (1.18)
Moments of $h_{r}^{-1}S_{r}$ also converge to moments of $\mathrm{d}\sigma_{s}$. In particular,
$\chi^{\mathbb{T}_{r}}(\beta_{c}+sw_{r})=r^{d}\langle S_{r}^{2}\rangle_{\beta_{c}+sw_{r}}^{\mathbb{T}_{r}}\sim r^{d}h_{r}^{2}f(s)=b_{d}^{2}\,f(s)\times\begin{cases}(\log r)^{1/2}r^{2}&(d=4)\\ r^{d/2}&(d>4),\end{cases}$ (1.19)
$\lim_{r\to\infty}g^{\mathbb{T}_{r}}(\beta_{c}+sw_{r})=3-\frac{\int_{\mathbb{R}}x^{4}\mathrm{d}\sigma_{s}}{(\int_{\mathbb{R}}x^{2}\mathrm{d}\sigma_{s})^{2}}.$ (1.20)
For $x\in{\mathbb{T}_{r}}$, the plateau is asymptotically the susceptibility divided by the volume:
$\tau^{\mathbb{T}_{r}}_{\beta_{c}+sw_{r}}(0,x)-\tau_{\beta_{c}}(0,x)\sim h_{r}^{2}f(s)=b_{d}^{2}\,f(s)\times\begin{cases}(\log r)^{1/2}r^{-2}&(d=4)\\ r^{-d/2}&(d>4).\end{cases}$ (1.21)
As observed in [33, Section 1.5.3], the right-hand side of (1.20) is strictly monotone increasing, and has limits 00 and $2$ as $s\to-\infty$ and $s\to+\infty$ respectively. Thus, assuming (1.20), the renormalised coupling constant varies monotonically over its possible range $[0,2]$ as the window is traversed. At the infinite-volume critical point, (1.20) gives
$\lim_{r\to\infty}g^{\mathbb{T}_{r}}(\beta_{c})=3-4\left(\frac{\Gamma(5/4)}{\Gamma(3/4)}\right)^{2}=0.81156\ldots\qquad(d\geq 4).$ (1.22)
The results of [33, 37] further suggest the following conjecture for the Ising model with FBC in dimensions $d\geq 4$.
Conjecture 1.6 (Free boundary conditions).
There is an effective critical point $\beta_{c,r}$ for the Ising model with free boundary conditions on the box $\Lambda_{r}$, which obeys
$\beta_{c,r}=\beta_{c}+v_{r},\qquad v_{r}\asymp\begin{cases}(\log r)^{1/3}r^{-2}&(d=4)\\ r^{-2}&(d>4),\end{cases}$ (1.23)
such that the statements of Conjecture 1.5 holds as stated when $\beta_{c}$ is replaced by $\beta_{c,r}$. In particular, we conjecture that the plateau phenomenon also occurs for free boundary conditions at the effective critical point.
The shift scale $v_{r}$ exceeds the window scale $w_{r}$, so $\beta_{c}$ is not inside the conjectured scaling window for FBC. This is consistent with theorems proving a Gaussian limit at $\beta_{c}$ under FBC [2, 20, 4, 35, 33, 37].
Closely related conjectures for percolation are discussed in Appendix C. Our analysis is based on the random current representation for the Ising model, which gives a geometric representation for correlation functions with similarities to bond percolation. In view of this parallel with percolation, the following intuitive picture of the percolation plateau provides intuition more generally.
Origin of the plateau.
For spread-out independent Bernoulli bond percolation on the torus $\mathbb{T}_{r}$ in dimensions $d>6$, it is proved in [28] that, in a window of width $r^{-d/3}$ containing the infinite-volume critical point $p_{c}$, the two-point function behaves as $|x|^{-(d-2)}+r^{-2d/3}$ and the susceptibility (expected cluster size) behaves as $r^{d/3}$. Also, it has been proved for sufficiently high dimensions (presumably true for all $d>6$) that at $p_{c}$ the largest cluster on $\mathbb{T}_{r}$ is of order $r^{2d/3}$, whereas with FBC the largest cluster has size $r^{4}$; see [26, Theorem 13.5], [26, Theorem 13.22], [3]. For PBC, if we assume that two points 00 and $x$ in $\mathbb{T}_{r}$ are connected to a largest cluster (in particular connected to each other) with probability $(r^{2d/3}/r^{d})^{2}$, we find a plateau term of the correct order $r^{-2d/3}$. For FBC, a similar computation gives a probability $(r^{4}/r^{d})^{2}=r^{-(d-2)+(d-6)}$, which is negligible compared to $\tau_{p_{c}}(0,x)\asymp|x|^{-(d-2)}$. This rough computation is consistent with a plateau for PBC and no plateau for FBC.
The equality of powers $r^{2d/3}$ and $r^{4}$ when $d=6$ indicates a special role for dimension $6$. No plateau is predicted for $d<6$, or indeed for other models such as spin systems below their upper critical dimension. For percolation in dimensions below $6$, hyperscaling plays a role [22]. Hyperscaling reflects the dimension-dependent manner in which critical clusters are constrained to fit into space. The distinction between a plateau above the critical dimension, and no plateau below the critical dimension, may be understood intuitively as due to the fact that in high dimensions a critical cluster that “wraps around” discovers new vertices, whereas in low dimensions the wrapping mostly encounters vertices already present in the cluster without wrapping. In [14], the role of wrapping above the upper critical dimension is studied numerically in related settings.
1.4 Structure of proof
We now state five propositions and prove our main results subject to these propositions. The proof of Theorem 1.1 mirrors the general strategy used for self-avoiding walk in [40, 30] and for percolation in [28]. However, there are model-specific elements in the proof of these propositions. For the Ising model, we will use (1.6)–(1.7) as well as the random current representation [2, 21].
We observe that $\tau_{\beta}$ and $\tau_{\beta}^{\mathbb{T}_{r}}$ are translation invariant (but $\tau_{\beta}^{\Lambda_{r}}$ is not), so we write $\tau_{\beta}(x)=\tau_{\beta}(0,x)$ and $\tau_{\beta}^{\mathbb{T}_{r}}(x)=\tau_{\beta}^{\mathbb{T}_{r}}(0,x)$ from now on. For $x\in{\mathbb{Z}^{d}}$, we define
$T_{\beta}^{(r)}(x)=\sum_{u\in\mathbb{Z}^{d}}\tau_{\beta}(x+ru).$ (1.24)
Proposition 1.7.
Let $d\geq 1$, $r\geq 3$, and $x\in{\mathbb{T}_{r}}$. Then $\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq T_{\beta}^{(r)}(x).$
Proposition 1.7 suggests decomposing $\tau_{\beta}^{\mathbb{T}_{r}}(x)$ as
$\tau_{\beta}^{\mathbb{T}_{r}}(x)=T_{\beta}^{(r)}(x)-\big{(}T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\big{)}.$ (1.25)
We estimate the two terms on the right-hand side separately as follows. Recall that $\chi(\beta)$ denotes the ${\mathbb{Z}^{d}}$ susceptibility.
Proposition 1.8.
Let $d>4$. There is a constant $c_{1}>0$ such that for all $r\geq 3$ and $\beta<\beta_{c}$,
$T_{\beta}^{(r)}(x)\leq\tau_{\beta}(x)+c_{1}\frac{{\chi(\beta)}}{r^{d}}\qquad(x\in{\mathbb{T}_{r}}).$ (1.26)
Proposition 1.9.
Let $d>4$ and $c_{3}=\frac{4}{9}c_{0}^{2}$, where $c_{0}$ is the constant of (1.7). There is a constant $c_{T}>0$ such that for all $r\geq 3$ and all $\beta\in[\beta_{c}-c_{3}r^{-2},\beta_{c})$,
$T_{\beta}^{(r)}(x)\geq\tau_{\beta}(x)+c_{T}\frac{{\chi(\beta)}}{r^{d}}\qquad(x\in{\mathbb{T}_{r}}).$ (1.27)
Proposition 1.10.
Let $d>4$. Let $c_{T}$ be the constant of Proposition 1.9. There are constants $c_{4},c_{5}>0$ such that for all $r\geq 3$ and $\beta\leq\beta_{c}-c_{4}r^{-d/2}$,
$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq\frac{1}{2}c_{T}\frac{\chi(\beta)}{r^{d}}\qquad(x\in{\mathbb{T}_{r}},\;\lVert x\rVert_{\infty}\geq c_{5}).$ (1.28)
Proof of Theorem 1.1.
The upper bound follows immediately from Propositions 1.7 and 1.8. Recall that $\beta_{*}=\beta_{c}-c_{4}r^{-d/2}$. For the lower bound, let $\beta\in[\beta_{c}-c_{3}r^{-2},\beta_{*}]$ and $\lVert x\rVert_{\infty}\geq c_{5}$. Then by (1.25), and by Propositions 1.9 and 1.10,
$\tau_{\beta}^{\mathbb{T}_{r}}(x)\geq\tau_{\beta}(x)+\left(1-\frac{1}{2}\right)c_{T}\frac{{\chi(\beta)}}{r^{d}}.$ (1.29)
This gives the desired result with $c_{2}=\frac{1}{2}c_{T}$. ∎
Proof of Corollary 1.2.
The upper bound follows from (1.9), $\tau_{\beta}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ by (1.6), and the fact that $\chi(\beta_{*})\asymp r^{-d/2}$ from (1.5).
For the lower bound, when $\lVert x\rVert_{\infty}\geq c_{5}$, we first observe that the lower bound (1.7) applies at $\beta=\beta_{*}$ provided that $r/2\leq c_{0}(\beta_{c}-\beta^{*})^{-1/2}$, i.e. $r^{(d-4)/4}\geq(2c_{0})^{-1}c_{4}^{1/2}$. It then follows from (1.10) and $\chi(\beta)\geq\chi(\beta_{*})\gtrsim r^{d/2}$ that
$\tau_{\beta}^{\mathbb{T}_{r}}(x)\geq\tau_{\beta_{*}}(x)+c_{2}\frac{\chi(\beta_{*})}{r^{d}}\gtrsim|\!|\!|x|\!|\!|^{-(d-2)}+r^{-d/2},$ (1.30)
as desired. When $\lVert x\rVert_{\infty}\leq c_{5}$, we let $m_{5}=\min_{\lVert x\rVert_{\infty}\leq c_{5}}\tau_{\beta_{c}/2}^{\Lambda_{2c_{5}}}(0,x)>0$. By the second Griffiths inequality, when $r$ is large we have
$\tau_{\beta_{*}}^{\mathbb{T}_{r}}(x)\geq\tau^{\Lambda_{2c_{5}}}_{\beta_{*}}(0,x)\geq m_{5}\geq\frac{m_{5}}{2}\frac{1}{|\!|\!|x|\!|\!|^{d-2}}+\frac{m_{5}}{2}\frac{1}{r^{d/2}},$ (1.31)
which completes the proof. ∎
The proof of Theorem 1.4 uses the following proposition, which we prove in Appendix B. The torus Ursell function $U_{4,\beta}^{\mathbb{T}_{r}}$ is defined by
$U_{4,\beta}^{\mathbb{T}_{r}}(0,x,y,z)=\langle\sigma_{0}\sigma_{x}\sigma_{y}\sigma_{z}\rangle_{\beta}^{\mathbb{T}_{r}}-\langle\sigma_{0}\sigma_{x}\rangle_{\beta}^{\mathbb{T}_{r}}\langle\sigma_{y}\sigma_{z}\rangle_{\beta}^{\mathbb{T}_{r}}-\langle\sigma_{0}\sigma_{y}\rangle_{\beta}^{\mathbb{T}_{r}}\langle\sigma_{x}\sigma_{z}\rangle_{\beta}^{\mathbb{T}_{r}}-\langle\sigma_{0}\sigma_{z}\rangle_{\beta}^{\mathbb{T}_{r}}\langle\sigma_{x}\sigma_{y}\rangle_{\beta}^{\mathbb{T}_{r}}.$ (1.32)
By the Lebowitz and Griffiths inequalities, $U_{4,\beta}^{\mathbb{T}_{r}}(0,x,y,z)\in[-2,0]$. The torus bubble diagram is defined by
${\sf B}(\beta)=\sum_{x\in\mathbb{T}_{r}}\tau_{\beta}^{\mathbb{T}_{r}}(x)^{2}.$ (1.33)
Proposition 1.11.
Let $d\geq 1$. For all $\beta>0$ and for all $r\geq 3$,
$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4,\beta}^{\mathbb{T}_{r}}(0,x,y,z)|\geq\frac{1}{128d^{2}}\cdot\min\Big{\{}\beta(\chi^{\mathbb{T}_{r}})^{2}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta},\frac{1}{{\sf B}}\Big{(}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta}\Big{)}^{2}\Big{\}}.$ (1.34)
Proof of Theorem 1.4.
All quantities in this proof are torus quantities, so to simplify the notation we drop the label $\mathbb{T}_{r}$ everywhere. By a direct computation using $S_{r}=r^{-d}\sum_{x\in\mathbb{T}_{r}}\sigma_{x}$,
$g(\beta)=-\frac{\langle S_{r}^{4}\rangle_{\beta}-3(\langle S_{r}^{2}\rangle_{\beta})^{2}}{(\langle S_{r}^{2}\rangle_{\beta})^{2}}=\frac{1}{\chi(\beta)^{2}r^{d}}\sum_{x,y,z}|U_{4,\beta}(0,x,y,z)|.$ (1.35)
We use Proposition 1.11 to bound from below the sum on the right-hand side. A lower bound on the derivative of the susceptibility, essentially due to [7] (see [19, (12.50)]), states that
$\beta\frac{\partial\chi}{\partial\beta}\geq\frac{2d\chi^{2}}{1+(2d)^{2}{\sf B}}\left[1-\frac{4d{\sf B}}{\chi}-\frac{{\sf B}}{\chi^{2}}\right].$ (1.36)
It follows from the upper bound of Corollary 1.2 that ${\sf B}(\beta_{*})\lesssim 1$. Combined with $\chi(\beta_{*})\asymp r^{d/2}$ from Corollary 1.3, we get
$g(\beta_{*})\gtrsim\frac{1}{r^{d}}\chi(\beta_{*})^{2}\asymp 1,$ (1.37)
and the proof is complete. ∎
2 Proof of Propositions 1.71.10
We now prove Propositions 1.71.10. The proofs of Propositions 1.7 and 1.10 depend on new diagrammatic estimates that we state as Theorem 2.1. We begin with the proofs of Propositions 1.8 and 1.9, which are easier.
2.1 Proof of Propositions 1.8 and 1.9
These two propositions provide upper and lower bounds on
$T_{\beta}^{(r)}(x)=\sum_{u\in\mathbb{Z}^{d}}\tau_{\beta}(x+ru),$ (2.1)
respectively. It is useful to observe that, since $\lVert x\rVert_{\infty}\leq\frac{r}{2}$ for all $x\in{\mathbb{T}_{r}}$, for all $0\neq u\in\mathbb{Z}^{d}$ we have
$\lVert x+ru\rVert_{\infty}\leq\lVert ru\rVert_{\infty}+\frac{r}{2}\leq\frac{3}{2}r\lVert u\rVert_{\infty},$ (2.2)
$\lVert x+ru\rVert_{\infty}\geq\lVert ru\rVert_{\infty}-\frac{r}{2}\geq\frac{1}{2}r\lVert u\rVert_{\infty}.$ (2.3)
Proposition 1.8 follows simply from (2.3) and the upper bound (1.6) on $\tau_{\beta}$, and we defer its proof to Appendix A where it is included in the elementary Lemma A.2. We prove Proposition 1.9 here, using (2.2) and the lower bound (1.7) on $\tau_{\beta}$.
Proof of Proposition 1.9.
For the proof, we wish to apply the lower bound (1.7) to all $x+ru$ with $x\in\Lambda_{r}$ and $\|u\|_{\infty}\leq M$, with a well-chosen $M\geq 1$. Since (1.7) holds when $\|x+ru\|_{\infty}\leq c_{0}(\beta_{c}-\beta)^{-1/2}$, it suffices to have
$M\geq 1,\qquad\frac{r}{2}+rM\leq c_{0}(\beta_{c}-\beta)^{-1/2}.$ (2.4)
We try $M=M(\beta,r)=\mu c_{0}(\beta_{c}-\beta)^{-1/2}r^{-1}$. The two inequalities of (2.4) then become
$r^{2}(\beta_{c}-\beta)\leq\mu^{2}c_{0}^{2},\qquad r^{2}(\beta_{c}-\beta)\leq 4(1-\mu)^{2}c_{0}^{2}.$ (2.5)
The choice $\mu=2/3$ makes the two right-hand sides equal, and in that case both inequalities assert $r^{2}(\beta_{c}-\beta)\leq\frac{4}{9}c_{0}^{2}$. Thus, if we choose $c_{3}=\frac{4}{9}c_{0}^{2}$ and require $\beta\in[\beta_{c}-c_{3}r^{-2},\beta_{c})$, then by using (1.7) and (2.2) in the definition of $T_{\beta}^{(r)}(x)$ we obtain
$T_{\beta}^{(r)}(x)-\tau_{\beta}(x)\gtrsim\sum_{1\leq\lVert u\rVert_{\infty}\leq M}\frac{1}{\lVert x+ru\rVert_{\infty}^{d-2}}\geq\sum_{1\leq\lVert u\rVert_{\infty}\leq M}\frac{1}{(\frac{3}{2}r\lVert u\rVert_{\infty})^{d-2}}\asymp\frac{M^{2}}{r^{d-2}}.$ (2.6)
By the definition of $M$ and the fact that $\chi(\beta)\asymp(\beta_{c}-\beta)^{-1}$ (by (1.5)), the right-hand side is bounded from below by a multiple of $r^{-d}\chi(\beta)$. This completes the proof. ∎
2.2 Diagrammatic estimates
The proofs of Propositions 1.7 and 1.10 rely on the following theorem, whose proof is our main work. The proof of Theorem 2.1 is based on the random current representation and a new coupling of the Ising model on the torus and on $\mathbb{Z}^{d}$, and is deferred to Section 3. For the statement of Theorem 2.1, given $x,x^{\prime}\in\mathbb{Z}^{d}$ we write $x\cong x^{\prime}$ if $x-x^{\prime}\in r\mathbb{Z}^{d}$, and we write $\pi x$ for the unique point in ${\Lambda_{r}}$ (which we identify with ${\mathbb{T}_{r}}$) such that $\pi x\cong x$. The bound (2.7) repeats the statement of Proposition 1.7.
Theorem 2.1.
Let $d\geq 1$. For any $x\in{\mathbb{T}_{r}}$,
$\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq T_{\beta}^{(r)}(x),$ (2.7)
$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq\sum_{x^{\prime}\cong x}\sum_{y}\sum_{\begin{subarray}{c}y^{\prime}\cong y\\ y^{\prime}\neq y\end{subarray}}\sum_{z}\tau_{\beta}(z)\tau_{\beta}(y-z)\tau_{\beta}(y^{\prime}-z)\tau_{\beta}^{\mathbb{T}_{r}}(\pi y^{\prime}-\pi z)\tau_{\beta}(x^{\prime}-y),$ (2.8)
where $x^{\prime},y,y^{\prime},z$ range over ${\mathbb{Z}^{d}}$ with the indicated restrictions.
The following corollary, which gives an upper bound that simplifies the presentation of (2.8), will be important. We define the torus $\star$ convolution for functions $f,g:\mathbb{T}_{r}\to\mathbb{R}$ (periodic functions on $\mathbb{Z}^{d}$) by
$(f\star g)(x)=\sum_{y\in\mathbb{T}_{r}}f(y)g(x-y)\qquad(x\in\mathbb{T}_{r}).$ (2.9)
We also use the ${\mathbb{Z}^{d}}$ convolution $(f*g)(x)=\sum_{y\in\mathbb{Z}^{d}}f(y)g(x-y)$ for absolutely summable functions on $\mathbb{Z}^{d}$, so it is important to mark the difference between $\star$ and *.
Corollary 2.2.
Let $d\geq 1$. For any $x\in{\mathbb{T}_{r}}$,
$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq 2\Big{(}T_{\beta}^{(r)}\star\big{(}T_{\beta}^{(r)}\big{)}^{2}[T_{\beta}^{(r)}-(\tau_{\beta}\circ\pi)]\star T_{\beta}^{(r)}\Big{)}(x).$ (2.10)
Proof.
We use (2.7) and the periodicity of $T_{\beta}^{(r)}$ to get $\tau_{\beta}^{\mathbb{T}_{r}}(\pi y^{\prime}-\pi z)\leq T_{\beta}^{(r)}(y-z)$. We also rewrite the sums over $x^{\prime},y^{\prime}$ using the definition of $T_{\beta}^{(r)}$. These allow us to bound the right-hand side of (2.8) by
$\sum_{y}\sum_{z}\tau_{\beta}(z)\tau_{\beta}(y-z)\big{(}T_{\beta}^{(r)}-\tau_{\beta}\big{)}(y-z)T_{\beta}^{(r)}(y-z)T_{\beta}^{(r)}(x-y).$ (2.11)
Since every point in ${\mathbb{Z}^{d}}$ can be written uniquely in the form $\Lambda_{r}+r{\mathbb{Z}^{d}}$, we can write
$x=\check{x},\quad y=\check{y}+ru,\quad z=\check{z}+rv\qquad(\check{x},\check{y},\check{z}\in\Lambda_{r},\ u,v\in{\mathbb{Z}^{d}}),$ (2.12)
so the two sums become four sums over $\check{y},\check{z},u,v$. Since $T_{\beta}^{(r)}$ is periodic, we have $T_{\beta}^{(r)}(y-z)=T_{\beta}^{(r)}(\check{y}-\check{z})$ and $T_{\beta}^{(r)}(x-y)=T_{\beta}^{(r)}(\check{x}-\check{y})$, independent of $u$. Hence, the sum over $u$ gives the upper bound
$\sum_{u}\big{(}\tau_{\beta}T_{\beta}^{(r)}-\tau_{\beta}^{2}\big{)}\big{(}\check{y}-\check{z}+r(u-v)\big{)}\leq T_{\beta}^{(r)}(\pi(\check{y}-\check{z}))^{2}-\tau_{\beta}(\pi(\check{y}-\check{z}))^{2},$ (2.13)
where in the subtracted term we kept only one term in the sum as an upper bound. We further bound the above difference of squares by $2T_{\beta}^{(r)}[T_{\beta}^{(r)}-(\tau_{\beta}\circ\pi)]$. Finally, we complete the proof with the observation that $\sum_{v}\tau_{\beta}(\check{z}+rv)=T_{\beta}^{(r)}(\check{z})$. ∎
Remark 2.3.
It is interesting to compare the diagrammatic upper bound (2.10) with diagrams in related models. The counterpart of $T_{\beta}^{(r)}-\tau_{\beta}^{\mathbb{T}_{r}}$ is denoted by $\psi-\psi^{\mathbb{T}}$ in [40, 28, 30]. For weakly or strictly self-avoiding walk, the bounds [40, (6.36)] and [30, (4.20)] can be written as
$T\star[\delta\cdot(T-\tau\circ\pi)]\star T,$ (2.14)
with $\delta$ the Kronecker delta. For percolation, the bounds of [28, (4.35), (4.42)] can be combined into the upper bound
$T\star[T\cdot(T-\tau\circ\pi)]\star T.$ (2.15)
For Ising, we found the bound
$T\star[T^{2}\cdot(T-\tau\circ\pi)]\star T.$ (2.16)
The $T^{2}$ in the Ising bound, compared to the $T$ in percolation, is responsible for lowering the critical dimension from $6$ to $4$. The three upper bounds are depicted in Figure 1.
Figure 1: From left to right, diagrammatic representations of the upper bounds for self-avoiding walk in (2.14), for percolation in (2.15), and for the Ising model in (2.16). The dotted line corresponds to $T-\tau\circ\pi$, which produces a factor $\chi/r^{d}$ for each model.
Proof of Proposition 1.10.
It follows from Corollary 2.2 and Proposition 1.8 that
$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\lesssim\frac{\chi(\beta)}{r^{d}}\Big{(}T_{\beta}^{(r)}\star\big{(}T_{\beta}^{(r)}\big{)}^{2}\star T_{\beta}^{(r)}\Big{)}(x)\qquad(x\in\mathbb{T}_{r}).$ (2.17)
By Lemma A.2, there is a constant $C>0$ such that
$(T_{\beta}^{(r)}\star\big{(}T_{\beta}^{(r)}\big{)}^{2}\star T_{\beta}^{(r)})(x)$ $\leq(\tau_{\beta}*\tau_{\beta}^{2}*\tau_{\beta})(x)+C\frac{{\chi(\beta)}^{2}}{r^{d}}+C\bigg{(}\frac{{\chi(\beta)}^{2}}{r^{d}}\bigg{)}^{2}.$ (2.18)
It follows from $\tau_{\beta}(x)\leq\tau_{\beta_{c}}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ and a standard convolution estimate [23, Proposition 1.7] that $(\tau_{\beta}*\tau_{\beta}^{2}*\tau_{\beta})(x)\lesssim|\!|\!|x|\!|\!|^{-(d-4)}$ when $d>4$, so this term can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ with $c_{5}$ large enough, uniformly in $\beta\leq\beta_{c}$. Also, for $\beta\leq\beta_{c}-c_{4}r^{-d/2}$, by (1.5) we have $r^{-d/2}\chi(\beta)\lesssim c_{4}^{-1}$, and we can make this as small as desired by taking $c_{4}$ large enough. With (2.18), this shows that the multiplier of ${\chi(\beta)}/r^{d}$ in (2.17) can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ and $c_{4},c_{5}$ large. In particular, it can be made smaller than $\frac{1}{2}c_{T}$, which completes the proof. ∎
3 Coupling and switching: Proof of Theorem 2.1
In this section, we introduce a new coupling between Ising models on ${\mathbb{T}_{r}}$ and on large finite subsets of $\mathbb{Z}^{d}$. The coupling uses the random current representation (see [16] for an introduction) and a percolation exploration process. To derive useful bounds, we develop a new switching lemma between currents on $\mathbb{T}_{r}$ and on large finite subsets of $\mathbb{Z}^{d}$, in Lemma 3.4.
3.1 Random currents and switching lemma
Definition 3.1.
Let $\mathcal{G}=(V,E)$ be a finite graph. A current $\mathbf{n}$ on $\mathcal{G}$ is a function from $E$ to $\mathbb{N}_{0}=\{0,1,\ldots\}$. We write $\Omega_{\mathcal{G}}$ for the set of all currents on $\mathcal{G}$. The set of sources of $\mathbf{n}$, denoted by $\partial\mathbf{n}$, is defined as
$\partial\mathbf{n}=\Big{\{}x\in V\mid\sum_{y:\>xy\in E}\mathbf{n}_{xy}\textup{is odd}\Big{\}}.$ (3.1)
It is often useful to interpret a current $\mathbf{n}$ as a multigraph on $V$ with $\mathbf{n}_{xy}$ distinct edges connecting $x$ and $y$ for all $xy\in E$. A sourceless current then can be seen as the edge count of a multigraph consisting of a union of loops. Adding sources to $\mathbf{n}$ amounts to adding a collection of paths connecting the elements of $\partial\mathbf{n}$ pairwise. In particular, a current $\mathbf{n}$ with sources $\partial\mathbf{n}=\{u,v\}$ can be seen as the edge count of a multigraph consisting of a union of loops together with a path from $u$ to $v$.
For $A\subset V$, we write $\sigma_{A}=\prod_{x\in A}\sigma_{x}$. It is well-known [2, 16] that the correlation functions of the Ising model (defined in (1.2)) are related to currents via
$\left\langle\sigma_{A}\right\rangle_{\beta}^{\mathcal{G}}=\dfrac{\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n})}{\sum_{\partial\mathbf{n}=\varnothing}w_{\beta}(\mathbf{n})}=\frac{Z^{A}_{\mathcal{G},\beta}}{Z^{\varnothing}_{\mathcal{G},\beta}},$ (3.2)
where
$w_{\beta}(\mathbf{n})=\prod_{xy\in E}\dfrac{\beta^{\mathbf{n}_{xy}}}{\mathbf{n}_{xy}!},\qquad Z^{A}_{\mathcal{G},\beta}=\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n}).$ (3.3)
For a set of currents $\mathcal{E}\subset\Omega_{\mathcal{G}}$, we also define
$Z^{A}_{\mathcal{G},\beta}[\mathcal{E}]=\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n})\mathds{1}\{\mathbf{n}\in\mathcal{E}\}.$ (3.4)
A current $\mathbf{n}$ induces a bond percolation configuration on $\mathcal{G}$ defined by $(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in E}$. The connectivity properties of this percolation configuration are known to play a crucial role in the analysis of the underlying Ising model. This motivates the following notation. For a current $\mathbf{n}$ on $\mathcal{G}$ and $x,y\in V$, we write $x\xleftrightarrow{\mathbf{n}\>}y$ if there exists a sequence of points $x_{0}=x,x_{1},\ldots,x_{k}=y$ such that $x_{i}x_{i+1}\in E$ and $\mathbf{n}_{x_{i}x_{i+1}}>0$ for all $0\leq i<k$.
We will use the following minor extension of the switching lemma of [21, 2] (see also [16, Lemma 7]). Given two graphs $\mathcal{G},\mathcal{H}$, two vertex sets $A\subset\mathcal{G}$, $B\subset\mathcal{H}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}\times\Omega_{\mathcal{H}}$, we write
$Z^{A,B}_{\mathcal{G},\mathcal{H},\beta}[\mathcal{E}]=\sum_{\begin{subarray}{c}\partial\mathbf{n}_{1}=A\\ \partial\mathbf{n}_{2}=B\end{subarray}}w_{\beta}(\mathbf{n}_{1})w_{\beta}(\mathbf{n}_{2})\mathds{1}\{(\mathbf{n}_{1},\mathbf{n}_{2})\in\mathcal{E}\}.$ (3.5)
We often drop the subscript $\beta$ when it does not play a role. For two sets $S_{1},S_{2}$, we write $S_{1}\Delta S_{2}=(S_{1}\setminus S_{2})\cup(S_{2}\setminus S_{1})$ for their symmetric difference.
Lemma 3.2 (Switching lemma).
Let $\mathcal{G}=(V,E)$ and $\mathcal{H}=(V^{\prime},E^{\prime})$ be finite graphs with $\mathcal{H}\subset\mathcal{G}$. For a current $\mathbf{n}$ on $\mathcal{G}$, let $\mathbf{n}_{|\mathcal{H}}$ denote its restriction to edges of $\mathcal{H}$. Then, for any $A\subset V$, $B\subset V^{\prime}$, $x,y\in V^{\prime}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}$,
$Z^{A,B}_{\mathcal{G},\mathcal{H}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}$ $=Z^{A\Delta\{x,y\},B\Delta\{x,y\}}_{\mathcal{G},\mathcal{H}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}.$ (3.6)
Proof.
By definition, $w(\mathbf{n}_{1})w(\mathbf{n}_{2})=w(\mathbf{n}_{1}+\mathbf{n}_{2})\binom{\mathbf{n}_{1}+\mathbf{n}_{2}}{\mathbf{n}_{1}}$. By the change of variables $\mathbf{m}=\mathbf{n}_{1}+\mathbf{n}_{2}$ and $\mathbf{n}=\mathbf{n}_{2}$, the left-hand side of (3.6) can be written as
$\sum_{\partial\mathbf{m}=A\Delta B}w(\mathbf{m})\mathds{1}\{\mathbf{m}\in\mathcal{E}\}\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.7)
We view $\mathbf{m},\mathbf{n}$ as multigraphs $\mathcal{M},\mathcal{N}$, and write $\partial\mathcal{N}\subset V^{\prime}$ for the vertices with an odd number of incident edges in $\mathcal{N}$. From this viewpoint, the sum over $\mathbf{n}$ above exactly counts the number of subgraphs $\mathcal{N}$ of $\mathcal{M}$ with $\partial\mathcal{N}=B$.
On the event $x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y$, there is a path in $\mathcal{M}$ from $x$ to $y$ that uses only edges in $E^{\prime}$. We denote such a path by $\mathcal{K}$. The involution $\mathcal{N}\mapsto\mathcal{N}\Delta\mathcal{K}$ then defines a bijection between subgraphs with source set $B$ and source set $B\Delta\{x,y\}$. Thus,
$\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}=\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\Delta\{x,y\}\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.8)
Undoing the change of variables produces the right-hand side of (3.6). ∎
3.2 A percolation exploration
We now consider bond percolation and introduce an exploration of the cluster of the origin that is particularly well-suited for the periodic setup to which we apply it. We follow the strategy initially introduced in [10] and revisited in [36, 25], with some modifications. We fix $r\geq 3$, and let $\mathcal{G}=(V,E)$ be a finite or infinite subgraph of $\mathbb{Z}^{d}$. Let $\omega=(\omega_{e})_{e\in E}$ be a bond percolation configuration on $\mathcal{G}$. We will define an exploration $\mathcal{C}(0)$ of the cluster of 00 in $\omega$ that contains at most one representative for each vertex and edge of ${\mathbb{T}_{r}}$, under the equivalence that $x\cong x^{\prime}$ if $x-x^{\prime}\in r\mathbb{Z}^{d}$ (for $x,x^{\prime}\in\mathcal{G}$). A simple example of the exploration is depicted in Figure 2.
Figure 2: An illustration of the exploration algorithm with $r=3$ and $\mathcal{G}=\mathbb{Z}^{2}$. Dotted boxes represent copies of $\Lambda_{3}$. The bond percolation configuration is shown in black and grey; the black edges are explored. Vertices not explored by the algorithm are unfilled. The graph $\mathcal{C}(0)$ consists of the black vertices and edges.
We use the following colouring scheme for vertices and edges of $\mathcal{G}$:
  • Vertices have three possible colourings: red (being or about to be explored), black (fully explored), or grey (has a coloured representative). We write $\mathsf{R}$, $\mathsf{B}$, $\mathsf{G}$ for the corresponding evolving sets.
  • Edges have two possible colourings: white (explored and closed) or black (explored and open). We let $\mathfrak{W}$, $\mathfrak{B}$ be the corresponding evolving sets.
The construction of $\mathcal{C}(0)$ is as follows. We fix an ordering of the vertices in $\mathcal{G}$.
  1. Step 0
[…]
  2. Step 1
    Let $u$ be the earliest element of $\mathsf{R}$.
    1. $(i)$
      Let $v$ be the earliest vertex adjacent to $u$ such that $v$ is not coloured grey and $uv$ is not coloured.
    2. $(ii)$
      If $\omega_{uv}=0$, add the edge $uv$ to $\mathfrak{W}$.
    3. $(iii)$
      If $\omega_{uv}=1$, add $v$ to $\mathsf{R}$ (if it is not already in it) and all its representatives $v^{\prime}\cong v$ ($v^{\prime}\neq v)$ to $\mathsf{G}$, add $uv$ to $\mathfrak{B}$.
    4. $(iv)$
      Restart from $(i)$ until all the edges containing $u$ are either coloured or connected to a grey vertex. Then add $u$ to $\mathsf{B}$.
  3. Step 2
    Restart from Step 1 until $\mathsf{R}=\varnothing$.
At the end of the exploration, we obtain $\mathsf{B},\mathsf{G},\mathfrak{B},\mathfrak{W}$. Note that all edges that have been explored belong to $\mathfrak{W}\cup\mathfrak{B}$. We define $\mathcal{C}(0)$ to be the graph $(\mathsf{B},\mathfrak{B})$.
We note the following properties of $\mathcal{C}(0)$:
  • By construction, $\mathcal{C}(0)$ contains at most one representative of each vertex and edge, and $\mathcal{C}(0)$ is connected. Any path in $\mathcal{C}(0)$ has at most $r^{d}$ vertices, so has length at most $r^{d}-1$.
[…]
    $\{\mathfrak{B}=B,\>\mathfrak{W}=W\}=\{\omega_{e}=1\text{ for }e\in B,\ \omega_{e}=0\text{ for }e\in W\}.$ (3.9)
    The $\subset$ inclusion is by definition of the algorithm. On the other hand, if $(B,W)$ is a valid output of the algorithm and $\omega$ is an element of the right-hand side, then running the exploration on $\omega$ will produce $\mathfrak{B}=B,\>\mathfrak{W}=W$, so the $\supset$ inclusion follows.
3.3 A coupled cluster exploration
We now apply the exploration introduced in Section 3.2 to a well-chosen percolation configuration induced by random currents. This allows us to perform switching between a current on a large finite subgraph $\mathcal{G}\subset\mathbb{Z}^{d}$ and a current on $\mathbb{T}_{r}$.
Let $\mathbf{m}$ be a current on $\mathcal{G}$ and $\mathbf{n}$ be a current on ${\mathbb{T}_{r}}$. We write $\hat{\mathbf{n}}$ for the lift of $\mathbf{n}$ to $\mathcal{G}$, defined by
$\hat{\mathbf{n}}(e)=\mathbf{n}(\pi e)\qquad(e\in\mathcal{G}),$ (3.10)
where $\pi:{\mathbb{Z}^{d}}\to{\mathbb{T}_{r}}$ is the canonical projection map. The current $\mathbf{m}+\hat{\mathbf{n}}$ induces a percolation configuration $(\mathds{1}_{(\mathbf{m}+\hat{\mathbf{n}})_{e}>0})_{e\in E}$ on $\mathcal{G}$. We run the exploration of Section 3.2 on this configuration and produce $(\mathsf{B},\mathsf{G},\mathfrak{B},\mathfrak{W})$ and $\mathcal{C}(0)=(\mathsf{B},\mathfrak{B})$.
Lemma 3.3 (Properties of the exploration).
Let $d\geq 1$, $r\geq 3$, $x\in\mathbb{T}_{r}$, and fix finite $\mathcal{G}\supset\Lambda_{2r^{d}}$. Let $\mathbf{m}\in\Omega_{\mathcal{G}}$, and let $\mathbf{n}\in\Omega_{\mathbb{T}_{r}}$ with $\partial\mathbf{n}=\{0,x\}$. Then the exploration $(\mathsf{B},\mathsf{G},\mathfrak{B},\mathfrak{W})$ of $(\mathds{1}_{(\mathbf{m}+\hat{\mathbf{n}})_{e}>0})_{e\in E}$ satisfies the following properties:
  1. (i)
    $(\mathbf{m}+\hat{\mathbf{n}})_{e}=0$ for all $e\in\mathfrak{W}$.
  2. (ii)
    $\hat{\mathbf{n}}_{e}=0$ for every representative of $e\in\mathfrak{W}$.
  3. (iii)
    There exists exactly one representative of $x$, which we denote by $x^{\prime}$, such that $x^{\prime}\in\mathcal{C}(0)$.
Proof.
The first two points follow directly from the algorithm of the exploration and the definition of $\hat{\mathbf{n}}$. The uniqueness of $x^{\prime}$ in $(iii)$ has been noted already in Section 3.2.
For existence, since $\partial\mathbf{n}=\{0,x\}$, there is an open path in $\mathbf{n}$ from 0 to $x$, which we label as $(y_{0},y_{1},\dots,y_{k})$ with $y_{0}=0$, $y_{k}=x$. We claim that, for any $j\in[0,k]$, there exists $y_{j}^{\prime}\in\mathcal{G}$ such that $y_{j}^{\prime}\cong y_{j}$ and $y_{j}^{\prime}\in\mathcal{C}(0)$. Then, we can take $x^{\prime}=y_{k}^{\prime}$. We prove the claim by induction on $j$. The $j=0$ case is trivial as we can take $y_{0}^{\prime}=0$.
Assume the claim for some $j<k$. Observe that $y_{j}^{\prime}\in\mathcal{C}(0)\subset\Lambda_{2r^{d}-1}$, because a path joining 00 to $y_{j}^{\prime}$ has length at most $r^{d}-1$. Let $z=y_{j}^{\prime}+(y_{j+1}-y_{j})\in\Lambda_{2r^{d}}\subset\mathcal{G}$, so $z\cong y_{j+1}$. If $z\in\mathcal{C}(0)$, we take $y_{j+1}^{\prime}=z$. Otherwise, $z$ must be grey because $\hat{\mathbf{n}}_{y_{j}^{\prime}z}=\mathbf{n}_{y_{j}y_{j+1}}>0$. This only happens when another representative $z^{\prime}$ of $z$ is explored before $z$. In that case, we take $y_{j+1}^{\prime}=z^{\prime}$. This completes the inductive step and the proof. ∎
We use the exploration to switch sources between currents on $\mathcal{G}$ and ${\mathbb{T}_{r}}$.
Lemma 3.4.
For any $C\subset\mathcal{G}$, $A\subset{\mathbb{T}_{r}}$, and $z\in\mathcal{G}$,
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[z\in\mathcal{C}(0)]=Z^{C\Delta\{0,z\},A\Delta\{0,\pi z\}}_{\mathcal{G},\mathbb{T}_{r}}[z\in\mathcal{C}(0)].$ (3.11)
In particular, for $x\in{\mathbb{T}_{r}}$, $A=\{0,x\}$, and $z=x^{\prime}$ with $\pi z=x$,
$Z^{\varnothing,0x}_{\mathcal{G},\mathbb{T}_{r}}[x^{\prime}\in\mathcal{C}(0)]=Z^{0x^{\prime},\varnothing}_{\mathcal{G},\mathbb{T}_{r}}[x^{\prime}\in\mathcal{C}(0)].$ (3.12)
Proof.
By conditioning on $(\mathfrak{B},\mathfrak{W})$, we have
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[z\in\mathcal{C}(0)]=\sum_{B\ni z,W}Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[(\mathfrak{B},\mathfrak{W})=(B,W)],$ (3.13)
where $B\ni z$ means that $z$ is incident to some edge in $B$. Given an admissible pair $(B,W)$ (i.e., realisable by the exploration and such that $B\ni z$), we make the decomposition
$\mathbf{m}$ $=\mathbf{m}_{1}+\mathbf{m}_{2},\qquad$ $\mathbf{m}_{1}(e)=\mathbf{m}(e)\mathds{1}\{e\in B\cup W\},$ (3.14)
$\mathbf{n}$ $=\mathbf{n}_{1}+\mathbf{n}_{2},\qquad$ $~{}\mathbf{n}_{1}(e)=\mathbf{n}(e)\mathds{1}\{e\in\pi(B)\}.$ (3.15)
Then $w(\mathbf{m})=w(\mathbf{m}_{1})w(\mathbf{m}_{2})$ and $w(\mathbf{n})=w(\mathbf{n}_{1})w(\mathbf{n}_{2})$. The current $\mathbf{n}_{2}$ is supported on $(\pi(B)\cup\pi(W))^{c}$, because $\mathbf{n}=0$ on $\pi(W)$. Also, since no edges connect to grey vertices in the exploration, the projection map $\pi$ restricts to an isomorphism between the graphs $B$ and $\pi(B)$, whose inverse we denote by $\iota:\pi(B)\to B$. Let $\tilde{\mathbf{n}}_{1}$ denote the restriction of $\hat{\mathbf{n}}_{1}$ to $B$. Then
$\iota^{-1}(\tilde{\mathbf{n}}_{1})=\mathbf{n}_{1},\quad w(\tilde{\mathbf{n}}_{1})=w(\mathbf{n}_{1}),\quad\iota^{-1}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{1}.$ (3.16)
Let $\mathcal{E}_{(B,W)}$ denote the event $\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$, which is measurable in $\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1}$ as indicated at (3.9). By conditioning on $\mathbf{m}_{2}$, $\mathbf{n}_{2}$,
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}]=\sum_{\mathbf{m}_{2},\mathbf{n}_{2}}w(\mathbf{m}_{2})w(\mathbf{n}_{2})\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\mathbf{n}_{1}=\partial\mathbf{n}_{2}\Delta A\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\},$ (3.17)
where we have $0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z$ because $B\ni z$. The sum over currents $\mathbf{n}_{1}$ on $\pi(B)$ can be replaced by a sum over currents $\tilde{\mathbf{n}}_{1}$ on $B$, with the source constraint $\partial\tilde{\mathbf{n}}_{1}=\iota(\partial\mathbf{n}_{2}\Delta A)$. By Lemma 3.2 applied to the graphs $B\cup W$ and $B$,
$\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\tilde{\mathbf{n}}_{1}=\iota(\partial\mathbf{n}_{2}\Delta A)\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\}$
$\hskip 42.67912pt=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\Delta\{0,z\}\\ \partial\tilde{\mathbf{n}}_{1}=[\iota(\partial\mathbf{n}_{2}\Delta A)]\Delta\{0,z\}\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\}.$ (3.18)
Now, we use $\partial\mathbf{n}_{1}=\iota^{-1}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{2}\Delta A\Delta\{0,\pi z\}$ to convert back to a sum over $\mathbf{n}_{1}$. The result is
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}]=Z^{C\Delta\{0,z\},A\Delta\{0,\pi z\}}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}].$ (3.19)
Summation over all admissible pairs $(B,W)$ then yields the desired result. ∎
We can now prove the first bound of Theorem 2.1.
Proof of (2.7).
Let $\mathcal{G}={\Lambda_{R}}$ for some $R\geq 2r^{d}$. We introduce an auxiliary sourceless current $\mathbf{m}$ on ${\Lambda_{R}}$. Then (3.2) and Lemma 3.3(iii) give
$\tau_{\beta}^{\mathbb{T}_{r}}(x)=\frac{Z_{\mathbb{T}_{r}}^{0x}}{Z_{\mathbb{T}_{r}}^{\varnothing}}=\frac{Z_{{\Lambda_{R}},\mathbb{T}_{r}}^{\varnothing,0x}}{Z_{{\Lambda_{R}},\mathbb{T}_{r}}^{\varnothing,\varnothing}}=\frac{Z^{\varnothing,0x}_{{\Lambda_{R}},{\mathbb{T}_{r}}}\Big{[}\bigsqcup_{x^{\prime}\cong x}\{x^{\prime}\in\mathcal{C}(0)\}\Big{]}}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{\varnothing,0x}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[x^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}.$ (3.20)
Using (3.12) and $\{x^{\prime}\in\mathcal{C}(0)\}\subset\{0\xleftrightarrow{\mathbf{m}}x^{\prime}\}$, we get
$\tau_{\beta}^{\mathbb{T}_{r}}(x)=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[x^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}\leq\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime}).$ (3.21)
Then (2.7) follows from the fact that $\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime})\leq\tau_{\beta}(x^{\prime})$ by the second Griffiths inequality. ∎
3.4 Proof of (2.8)
Let $R\geq 2r^{d}$. We define
$T_{\beta}^{(r,R)}(x)=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime})\qquad(x\in\mathbb{T}_{r}).$ (3.22)
Then, by the first equality of (3.21),
$T_{\beta}^{(r,R)}(x)-\tau^{\mathbb{T}_{r}}_{\beta}(x)=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},\mathbb{T}_{r}}[x^{\prime}\notin\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},\mathbb{T}_{r}}}.$ (3.23)
The inequality (2.8) is an immediate consequence of the following lemma, by taking the limit $R\to\infty$.
Lemma 3.5.
Let $d\geq 1$ and $R\geq 2r^{d}$. For any $x\in{\mathbb{T}_{r}}$, we have
$T_{\beta}^{(r,R)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq\sum_{x^{\prime}\cong x}\sum_{y}\sum_{\begin{subarray}{c}y^{\prime}\cong y\\ y^{\prime}\neq y\end{subarray}}\sum_{z}\tau^{\Lambda_{R}}_{\beta}(0,z)\tau^{\Lambda_{R}}_{\beta}(z,y)\tau^{\Lambda_{R}}_{\beta}(z,y^{\prime})\tau^{\mathbb{T}_{r}}_{\beta}(\pi y^{\prime}-\pi z)\tau^{\Lambda_{R}}_{\beta}(y,x^{\prime}),$ (3.24)
where $x^{\prime},y,y^{\prime},z$ range over ${\Lambda_{R}}$ with the indicated restrictions.
The proof proceeds by bounding the terms on the right-hand side of (3.23). For each $x^{\prime}$, since $\partial\mathbf{m}=\{0,x^{\prime}\}$, there is an open path in $\mathbf{m}$ from 00 to $x^{\prime}$. When $x^{\prime}\not\in\mathcal{C}(0)$, every such path must contain a grey site. (Note that $x^{\prime}$ might be uncoloured, because the sources do not guarantee that there is an $x^{\prime\prime}\in\mathcal{C}(0)$ with $x^{\prime\prime}\cong x$.) We let $y\in{\Lambda_{R}}$ denote the last grey site on one of the paths, so that $y\xleftrightarrow{\mathbf{m}\>}x^{\prime}$ using only edges in $(\mathfrak{B}\cup\mathfrak{W})^{c}$. Also, since $y$ is grey, there exists another representative $y^{\prime}\cong y$, $y^{\prime}\neq y$ such that $y^{\prime}\in\mathcal{C}(0)$. This proves the inclusion
$\{x^{\prime}\not\in\mathcal{C}(0)\}\subset\bigcup_{y\in{\Lambda_{R}}}\bigcup_{\begin{subarray}{c}y^{\prime}\cong y\\ y^{\prime}\neq y\end{subarray}}\{y^{\prime}\in\mathcal{C}(0)\}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(\mathfrak{B}\cup\mathfrak{W})^{c}\}.$ (3.25)
In the following, we regard an edge subset (such as $(\mathfrak{B}\cup\mathfrak{W})^{c}$) as the graph whose edges are those in that subset and whose vertices are the vertices comprising those edges.
Lemma 3.6.
Let $\mathcal{A}(x^{\prime},y,y^{\prime})=\{y^{\prime}\in\mathcal{C}(0)\}\cap\{y\xleftrightarrow{\mathbf{m}}x^{\prime}\text{ in }(\mathfrak{B}\cup\mathfrak{W})^{c}\}$. Then,
$Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}(x^{\prime},y,y^{\prime})]\leq Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime}).$ (3.26)
Proof.
For simplicity, we write $\mathcal{A}=\mathcal{A}(x^{\prime},y,y^{\prime})$. Since the last event in (3.25) is a connection outside of $\mathfrak{B}\cup\mathfrak{W}$, we can use it to switch sources with a third independent current without affecting $\mathfrak{B}\cup\mathfrak{W}$, as follows. Let $\bm{\ell}$ be an independent current on ${\Lambda_{R}}$ with $\partial\bm{\ell}=\varnothing$. We again condition on the event $\mathcal{E}_{(B,W)}=\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$. We decompose
$\mathbf{m}=\mathbf{m}_{1}+\mathbf{m}_{2},\qquad\mathbf{m}_{1}(e)=\mathbf{m}(e)\mathds{1}\{e\in B\cup W\},$ (3.27)
and similarly for $\bm{\ell}=\bm{\ell}_{1}+\bm{\ell}_{2}$. By conditioning on $\mathbf{m}_{1}$ and $\bm{\ell}_{1}$,
$Z^{\varnothing}_{\Lambda_{R}}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]=Z^{\varnothing}_{\Lambda_{R}}\sum_{B,W}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}\cap\mathcal{E}_{(B,W)}]$
$\quad=\sum_{B\ni y^{\prime},W}Z^{\varnothing,0x^{\prime},\varnothing}_{{\Lambda_{R}},{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(B\cup W)^{c}\}]$
$\quad=\sum_{B\ni y^{\prime},W}\sum_{\begin{subarray}{c}\bm{\ell}_{1},\mathbf{m}_{1}\\ \partial\mathbf{n}=\varnothing\end{subarray}}w(\bm{\ell}_{1})w(\mathbf{m}_{1})w(\mathbf{n})\mathds{1}_{\mathcal{E}_{(B,W)}}\!\!\!\!\!\!\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}.$ (3.28)
Since $\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\subset\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}$, Lemma 3.2 applied with $\mathcal{G}=\mathcal{H}=(B\cup W)^{c}$ gives
$\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\leq\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\Delta\{y,x^{\prime}\}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,y\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}.$ (3.29)
We insert this into (3.4), and use $\mathds{1}\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}\leq 1$, to get
$Z^{\varnothing}_{\Lambda_{R}}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]\leq\sum_{B\ni y^{\prime},W}Z^{yx^{\prime},0y,\varnothing}_{{\Lambda_{R}},{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}]=Z^{yx^{\prime}}_{\Lambda_{R}}Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)].$ (3.30)
The claim then follows from division by $Z^{\varnothing}_{\Lambda_{R}}$ and (3.2). ∎
In the proof of Lemma 3.5, we use the notion of the backbone of a current [6, Section 4] (more on the backbone can be found in [5, 35]). A current $\mathbf{m}$ on a finite graph $\mathcal{G}$ with sources $\partial\mathbf{m}=\{u,v\}$ must have at least one path from $u$ to $v$ in the percolation configuration induced by $\mathbf{m}$. The backbone of $\mathbf{m}$ is an exploration of one of these paths. We fix some arbitrary ordering $\prec$ of the edges of $\mathcal{G}$.
Definition 3.7 (The backbone of a current).
Let $\mathbf{m}\in\Omega_{\mathcal{G}}$ with $\partial\mathbf{m}=\{u,v\}$. The backbone of $\mathbf{m}$, denoted $\Gamma_{\mathbf{m}}$, is the unique oriented and edge self-avoiding path from $u$ to $v$, supported on edges with $\mathbf{m}$ odd, that is minimal for $\prec$. The backbone $\Gamma_{\mathbf{m}}=\{x_{i}x_{i+1}:0\leq i<k\}$ is obtained via the following exploration process:
  1. $(i)$
    Let $x_{0}=u$. The first edge $x_{0}x_{1}$ of $\Gamma_{\mathbf{m}}$ is the earliest of all the edges $e$ emerging from $x_{0}$ with $\mathbf{m}_{e}$ odd. Thus, all edges $e\ni x_{0}$ satisfying $e\prec x_{0}x_{1}$ and $e\neq x_{0}x_{1}$ are explored and have $\mathbf{m}_{e}$ even.
  2. $(ii)$
    Inductively, each edge $x_{i}x_{i+1}$, $i\geq 1$, is the earliest of all edges $e$ emerging from $x_{i}$ that have not been explored previously, and for which $\mathbf{m}_{e}$ is odd.
  3. $(iii)$
    The exploration stops when it reaches a vertex from which no more odd, non-explored edges are available. This happens at $x_{k}=v$.
Let $\overline{\Gamma}_{\mathbf{m}}$ denote the set of explored edges, i.e. $\Gamma_{\mathbf{m}}$ together with all explored even edges.
By definition, $\overline{\Gamma}_{\mathbf{m}}$ is determined by $\Gamma_{\mathbf{m}}$ and the ordering $\prec$, and similarly $\Gamma_{\mathbf{m}}$ can be determined from $\overline{\Gamma}_{\mathbf{m}}$ by selecting odd edges in $\overline{\Gamma}_{\mathbf{m}}$ according to $\prec$. Also, the current $\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}$, defined to be the restriction of $\mathbf{m}$ to the complement $\overline{\Gamma}_{\mathbf{m}}^{c}$, is sourceless.
Proof of Lemma 3.5.
By (3.23), (3.25), a union bound, and Lemma 3.6,
$T_{\beta}^{(r,R)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq\sum_{x^{\prime}\cong x}\sum_{y}\sum_{\begin{subarray}{c}y^{\prime}\cong y\\ y^{\prime}\neq y\end{subarray}}\frac{Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime}).$ (3.31)
The connections used to bound the above right-hand side are depicted in Figure 3. We first apply Lemma 3.4 to switch on the event $y^{\prime}\in\mathcal{C}(0)$. Using Lemma 3.4 with $z=y^{\prime}$, $C=\{0,y\}$, and $A=\varnothing$, we get
$Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]=Z^{yy^{\prime},0\pi y^{\prime}}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)].$ (3.32)
Figure 3: An illustration of the connections relevant to the right-hand side of (3.31). The region delimited by the dotted line corresponds to the exploration $\mathcal{C}(0)$. The connection from $y$ to $x^{\prime}$ (in green) does not involve any edge of $(\mathfrak{B},\mathfrak{W})$. The point $z$ is defined by (3.33). Following the notation below (3.33): the red path corresponds to $\Gamma_{\mathbf{m}}$ after the switching performed in (3.32), the blue path corresponds to a connection in $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$. The additional torus two-point function appearing in the right-hand side of (3.24) is produced by the switching of (3.4).
Since $\partial\mathbf{m}=\{y,y^{\prime}\}$ now, we can form the backbone $\Gamma_{\mathbf{m}}$ of $\mathbf{m}$ from $y^{\prime}$ to $y$. Let $\overline{\Gamma}_{\mathbf{m}}$ be the set of edges explored during the formation of $\Gamma_{\mathbf{m}}$. Let $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$ denote the result of the exploration process of Section 3.2 performed on the current $\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}$ (on the graph ${\Lambda_{R}}\setminus\overline{\Gamma}_{\mathbf{m}}$). We claim that
$\{y^{\prime}\in\mathcal{C}(0)\}\subset\bigcup_{z\in{\Lambda_{R}}}\{z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}\>}y^{\prime}\}\cap\{z\in\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)\}.$ (3.33)
Indeed, if $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)=\mathcal{C}(0)$, then (3.33) holds with $z=y^{\prime}$. If $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)\neq\mathcal{C}(0)$, then there is a first edge $zv\in\overline{\Gamma}_{\mathbf{m}}$ used in the construction of $\mathcal{C}(0)$ that makes the two exploration processes different. For this edge, we must have $z\in\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$ and $\mathbf{m}_{zv}>0$. Since one of $z,v$ must be in $\Gamma_{\mathbf{m}}$, we see that $z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}}y^{\prime}$ also. This proves (3.33).
Conditional on $\Gamma_{\mathbf{m}}=\gamma$, we set $\bar{\gamma}=\overline{\Gamma}_{\mathbf{m}}$ and decompose $\mathbf{m}$ as
$\mathbf{m}=\mathbf{k}+\bm{\ell},\qquad\mathbf{k}(e)=\mathbf{m}(e)\mathds{1}\{e\in\bar{\gamma}\},$ (3.34)
so $\mathbf{k},\bm{\ell}$ are currents on $\bar{\gamma}$, ${\Lambda_{R}}\setminus\bar{\gamma}$ respectively. Note that $w(\mathbf{m})=w(\mathbf{k})w(\bm{\ell})$, $\partial\mathbf{k}=\{y,y^{\prime}\}$, $\partial\bm{\ell}=\varnothing$, and $\Gamma_{\mathbf{m}}=\Gamma_{\mathbf{k}}$. Using (3.33),
$Z^{yy^{\prime},0\pi y^{\prime}}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]\leq\sum_{\begin{subarray}{c}z\in{\Lambda_{R}}\end{subarray}}\sum_{\gamma}\sum_{\partial\mathbf{k}=\{y,y^{\prime}\}}w(\mathbf{k})\mathds{1}\{\Gamma_{\mathbf{k}}=\gamma\}\mathds{1}\{z\xleftrightarrow{\overline{\Gamma}_{\mathbf{k}}\>}y^{\prime}\}Z^{\varnothing,0\pi y^{\prime}}_{{\Lambda_{R}}\setminus\bar{\gamma},\mathbb{T}_{r}}[z\in\mathcal{C}_{\bm{\ell}+\hat{\mathbf{n}}}(0)].$ (3.35)
By Lemma 3.4 with $\mathcal{G}={\Lambda_{R}}\setminus\bar{\gamma}$,
$Z^{\varnothing,0\pi y^{\prime}}_{{\Lambda_{R}}\setminus\bar{\gamma},\mathbb{T}_{r}}[z\in\mathcal{C}_{\bm{\ell}+\hat{\mathbf{n}}}(0)]$ $=Z^{0z,\pi z\pi y^{\prime}}_{{\Lambda_{R}}\setminus\bar{\gamma},\mathbb{T}_{r}}[z\in\mathcal{C}_{\bm{\ell}+\hat{\mathbf{n}}}(0)]$
$\leq Z_{{\Lambda_{R}}\setminus\bar{\gamma}}^{0z}Z^{\pi z\pi y^{\prime}}_{\mathbb{T}_{r}}=Z_{{\Lambda_{R}}\setminus\bar{\gamma}}^{\varnothing}\tau_{\beta}^{{\Lambda_{R}}\setminus\bar{\gamma}}(0,z)Z^{\pi z\pi y^{\prime}}_{\mathbb{T}_{r}}.$ (3.36)
Then we use $\tau_{\beta}^{{\Lambda_{R}}\setminus\bar{\gamma}}\leq\tau_{\beta}^{\Lambda_{R}}$ to obtain
$Z^{yy^{\prime},0\pi y^{\prime}}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]$ $\leq\sum_{\begin{subarray}{c}z\in{\Lambda_{R}}\end{subarray}}Z^{\pi z\pi y^{\prime}}_{\mathbb{T}_{r}}\sum_{\gamma}\sum_{\partial\mathbf{k}=\{y,y^{\prime}\}}w(\mathbf{k})\mathds{1}\{\Gamma_{\mathbf{k}}=\gamma\}\mathds{1}\{z\xleftrightarrow{\overline{\Gamma}_{\mathbf{k}}\>}y^{\prime}\}Z_{{\Lambda_{R}}\setminus\bar{\gamma}}^{\varnothing}\tau_{\beta}^{{\Lambda_{R}}}(0,z)$
$=\sum_{\begin{subarray}{c}z\in{\Lambda_{R}}\end{subarray}}Z^{\pi z\pi y^{\prime}}_{\mathbb{T}_{r}}\sum_{\gamma}Z^{yy^{\prime}}_{\Lambda_{R}}[\Gamma_{\mathbf{m}}=\gamma,z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}\>}y^{\prime}]\tau_{\beta}^{{\Lambda_{R}}}(0,z)$
$=\sum_{\begin{subarray}{c}z\in{\Lambda_{R}}\end{subarray}}Z^{\pi z\pi y^{\prime}}_{\mathbb{T}_{r}}Z^{yy^{\prime}}_{\Lambda_{R}}[z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}\>}y^{\prime}]\tau_{\beta}^{{\Lambda_{R}}}(0,z).$ (3.37)
For the middle factor on the right-hand side, we introduce an auxiliary sourceless current $\mathbf{m}^{\prime}$ on ${\Lambda_{R}}$, and use Lemma 3.2 with $\mathcal{G}=\mathcal{H}={\Lambda_{R}}$:
$Z^{yy^{\prime}}_{\Lambda_{R}}[z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}\>}y^{\prime}]\leq\frac{Z^{yy^{\prime},\varnothing}_{{\Lambda_{R}},{\Lambda_{R}}}[z\xleftrightarrow{\mathbf{m}+\mathbf{m}^{\prime}}y^{\prime}]}{Z^{\varnothing}_{{\Lambda_{R}}}}=\frac{Z^{yz,zy^{\prime}}_{{\Lambda_{R}},{\Lambda_{R}}}}{Z^{\varnothing}_{{\Lambda_{R}}}}=Z^{yz}_{{\Lambda_{R}}}\tau_{\beta}^{{\Lambda_{R}}}(z,y^{\prime}).$ (3.38)
Inserting this yields
$\frac{Z^{yy^{\prime},0\pi y^{\prime}}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}$ $\leq\sum_{\begin{subarray}{c}z\in{\Lambda_{R}}\end{subarray}}\frac{Z^{\pi z\pi y^{\prime}}_{\mathbb{T}_{r}}}{Z^{\varnothing}_{\mathbb{T}_{r}}}\frac{Z^{yz}_{{\Lambda_{R}}}}{Z^{\varnothing}_{\Lambda_{R}}}\tau_{\beta}^{{\Lambda_{R}}}(z,y^{\prime})\tau_{\beta}^{{\Lambda_{R}}}(0,z)$
$=\sum_{\begin{subarray}{c}z\in{\Lambda_{R}}\end{subarray}}\tau_{\beta}^{\mathbb{T}_{r}}(\pi y^{\prime}-\pi z)\tau_{\beta}^{\Lambda_{R}}(z,y)\tau_{\beta}^{{\Lambda_{R}}}(z,y^{\prime})\tau_{\beta}^{{\Lambda_{R}}}(0,z).$ (3.39)
With (3.31) and (3.32), this concludes the proof. ∎
Appendix A Torus convolution estimates
We present two elementary convolution lemmas. The first is a special case of [34, Lemma 3.4]. We include the simple proof for completeness, and as an opportunity to correct an omission in the proof of [34, Lemma 3.4] (repeated in [33, Lemma 4.2]) which however does not change the conclusion of those lemmas.
Lemma A.1.
Let $d\geq 1$, $r\geq 1$, $a>0$ and $\mu>0$. There is a constant $C=C(d,a)>0$ such that
$\sum\limits_{u\in\mathbb{Z}^{d}:u\neq 0}\frac{1}{\|x+ru\|_{\infty}^{d-a}}e^{-\mu\|x+ru\|_{\infty}}$ $\leq Ce^{-\frac{1}{4}\mu r}\frac{1}{\mu^{a}r^{d}}\qquad(x\in\Lambda_{r}).$ (A.1)
Proof.
For all nonzero $u\in\mathbb{Z}^{d}$ and $r\geq 1$, we have $\lVert x+ru\rVert_{\infty}\geq\frac{1}{2}\|ru\|_{\infty}$ by (2.3). Hence,
$\sum_{u\in\mathbb{Z}^{d}:u\neq 0}\frac{1}{\|x+ru\|_{\infty}^{d-a}}e^{-\mu\|x+ru\|_{\infty}}$ $\leq 2^{d-a}\sum_{N=1}^{\infty}\sum_{u:\|u\|_{\infty}=N}\frac{1}{\|ru\|_{\infty}^{d-a}}e^{-\frac{1}{2}\mu\|ru\|_{\infty}}$
$\lesssim r^{a-d}e^{-\frac{1}{4}\mu r}\sum_{N=1}^{\infty}N^{a-1}e^{-\frac{1}{4}\mu rN}.$ (A.2)
We set $f(x)=x^{a-1}e^{-bx}$ for $x>0$, and $b=\frac{1}{4}\mu r>0$, so the goal is to prove that $\sum_{N=1}^{\infty}f(N)\lesssim b^{-a}$. If $a\in(0,1]$ then $f$ is a decreasing function, so the sum is bounded by $\int_{0}^{\infty}f(x)\mathrm{d}x=b^{-a}\Gamma(a)$. If $a>1$ then $f(x)$ is increasing for $x<x_{0}=\frac{a-1}{b}$ and is decreasing for $x>x_{0}$. We set $N_{0}=\lfloor x_{0}\rfloor$ and bound the sum by
$\bigg{(}\int_{1}^{N_{0}}f(x)\mathrm{d}x+f(N_{0})\bigg{)}\mathds{1}_{N_{0}\geq 1}+f(N_{0}+1)+\int_{N_{0}+1}^{\infty}f(x)\mathrm{d}x.$ (A.3)
The sum of the two integrals is again bounded by $\int_{0}^{\infty}f(x)\mathrm{d}x=b^{-a}\Gamma(a)$. For the other terms, if $N_{0}=0$ we use that $f(N_{0}+1)=f(1)=e^{-b}\leq c_{a}b^{-a}$ for all $b>0$. If $N_{0}\geq 1$, we use
$f(N_{0}),f(N_{0}+1)\leq\max_{x\geq 0}f(x)=\frac{1}{b^{a-1}}\max_{t\geq 0}t^{a-1}e^{-t}\leq\frac{a-1}{b^{a}}c_{a},$ (A.4)
where the last inequality is due to $b\leq a-1$ since $x_{0}\geq N_{0}\geq 1$. This completes the proof. ∎
The next lemma is a minor extension of [34, Lemma 3.6]. The definition of $T_{\beta}^{(r)}(x)$ is in (1.24), $\star$ denotes convolution on $\mathbb{T}_{r}$ and * denotes convolution on $\mathbb{Z}^{d}$.
Lemma A.2.
Let $d>4$. There is a constant $C>0$ such that for all $r\geq 3$, $\beta<\beta_{c}$, and $x\in{\mathbb{T}_{r}}$,
$T_{\beta}^{(r)}(x)$ $\leq\tau_{\beta}(x)+C\frac{\chi(\beta)}{r^{d}},$ (A.5)
$(T_{\beta}^{(r)}\star T_{\beta}^{(r)})(x)$ $\leq(\tau_{\beta}*\tau_{\beta})(x)+C\frac{\chi(\beta)^{2}}{r^{d}},$ (A.6)
$(\big{(}T_{\beta}^{(r)}\big{)}^{2}\star T_{\beta}^{(r)}\star T_{\beta}^{(r)})(x)$ $\leq(\tau_{\beta}^{2}*\tau_{\beta}*\tau_{\beta})(x)+C\frac{{\chi(\beta)}^{2}}{r^{d}}+C\bigg{(}\frac{{\chi(\beta)}^{2}}{r^{d}}\bigg{)}^{2}.$ (A.7)
Proof.
We suppress the subscript $\beta$ and superscript $r$ in the notation, and write $m=(\beta_{c}-\beta)^{1/2}$. Then $m^{-2}\asymp\chi$ by (1.5). For (A.5), we separate the $u=0$ term in the definition (1.24) of $T(x)$, and then use (1.6) and the $a=2$ case of Lemma A.1, to obtain
$T(x)$ $\leq\tau(x)+\sum_{u\neq 0}\frac{C_{0}}{\|x+ru\|_{\infty}^{d-2}}e^{-c_{0}m\|x+ru\|_{\infty}}\leq\tau(x)+C\frac{\chi}{r^{d}}.$ (A.8)
For (A.6), we first observe that
$(T\star T)(x)$ $=\sum_{y\in{\mathbb{T}_{r}}}\sum_{u,v\in\mathbb{Z}^{d}}\tau(y+ru)\tau(x-y+rv)$
$=\sum_{w\in\mathbb{Z}^{d}}\sum_{y\in{\mathbb{T}_{r}}}\sum_{v\in\mathbb{Z}^{d}}\tau(y-rv+rw)\tau(x-y+rv)=\sum_{w\in\mathbb{Z}^{d}}(\tau*\tau)(x-rw),$ (A.9)
where in the second equality we replaced $u$ by $w-v$, and in the third we observe that $y-rv$ ranges over $\mathbb{Z}^{d}$ under the indicated summations. The $w=0$ term is $(\tau*\tau)(x)$. For the remaining terms, we use the convolution estimate of [23, Proposition 1.7] and (1.6) to get $(\tau*\tau)(z)\lesssim|\!|\!|z|\!|\!|^{-(d-4)}\exp[-c_{0}m\|z\|_{\infty}]$. We then apply Lemma A.1 with $a=4$ to see that
$\sum_{w\neq 0}(\tau*\tau)(x-rw)\lesssim\frac{\chi^{2}}{r^{d}},$ (A.10)
which gives (A.6).
For (A.7), by applying the inequality $T\leq\tau\circ\pi+C\chi/r^{d}$ from (A.5), we get three terms by expanding the $T^{2}$ in $(T^{2}\star T\star T)(x)$. The first term is
$((\tau\circ\pi)^{2}\star T\star T)(x)$ $=\sum_{y,z\in\mathbb{T}_{r}}\sum_{v,w\in\mathbb{Z}^{d}}\tau^{2}(y)\tau(z+rw-y-rv)\tau(x-z-rw)$
$=\sum_{y\in\mathbb{T}_{r}}\sum_{v\in\mathbb{Z}^{d}}\tau^{2}(y)(\tau*\tau)(x-y-rv)$
$\leq\sum_{u\in{\mathbb{Z}^{d}}}(\tau^{2}*\tau*\tau)(x+ru).$ (A.11)
Since $d>4$, by [23, Proposition 1.7], $\tau^{2}*\tau*\tau$ obeys the same upper bound as the bound used on $\tau*\tau$ in the proof of (A.6), so we again have a bound $\chi^{2}/r^{d}$ for the sum over nonzero $u$. The cross term is
$2C\frac{\chi}{r^{d}}((\tau\circ\pi)\star T\star T)(x)$ $\lesssim\frac{\chi}{r^{d}}\lVert\tau\rVert_{L^{1}({\mathbb{T}_{r}})}\lVert T\star T\rVert_{\infty}\lesssim\frac{\chi}{r^{d}}\chi\Big{(}1+\frac{\chi^{2}}{r^{d}}\Big{)},$ (A.12)
by (A.6) together with the fact that $\|\tau*\tau\|_{\infty}\lesssim 1$ when $d>4$. The last term is
$\Big{(}C\frac{\chi}{r^{d}}\Big{)}^{2}(1\star T\star T)(x)\lesssim\frac{\chi^{2}}{r^{2d}}\lVert 1\rVert_{\infty}\lVert T\rVert_{L^{1}({\mathbb{T}_{r}})}^{2}=\frac{\chi^{2}}{r^{2d}}\chi^{2},$ (A.13)
since $\lVert T\rVert_{L^{1}({\mathbb{T}_{r}})}=\chi$. This completes the proof. ∎
Appendix B Proof of Proposition 1.11
We now prove Proposition 1.11. Although this differential inequality is essentially the same as [19, (12.118)], the justification in [19] involves an unpublished work of Fernández, and to our knowledge a proof has not been published elsewhere. We therefore present a proof here. We work on the torus and to simplify the notation we drop the labels $\beta$ and $\mathbb{T}_{r}$. We apply the dilution trick, following the strategy of [6].
We begin with a reduction to the following proposition. Although we only need the nearest-neighbour case, we prove the result for a general Hamiltonian
$H(\sigma)=-\!\!\!\!\sum_{\{x,y\}\subset{\mathbb{T}_{r}}}J_{x,y}\sigma_{x}\sigma_{y},$ (B.1)
[…]
$w(\mathbf{n})=\prod_{\{x,y\}\subset{\mathbb{T}_{r}}}\dfrac{(\beta J_{x,y})^{\mathbf{n}_{xy}}}{\mathbf{n}_{xy}!}.$ (B.2)
Recall that the torus bubble diagram is ${\sf B}(\beta)=\sum_{x\in\mathbb{T}_{r}}\tau_{\beta}(x)^{2}$.
Proposition B.1.
Let $d\geq 1$. For all $\beta>0$, for all $r\geq 3$, and for all $p\in[0,1]$,
$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4}(0,x,y,z)|\geq 2p\chi^{2}\beta\frac{\partial\chi}{\partial\beta}-32p^{2}\beta^{2}|J|^{2}\chi^{4}{\sf B}.$ (B.3)
Proof of Proposition 1.11.
The lower bound in (B.3) can be written as
$ap-bp^{2}\quad\text{with}\quad a=2\chi^{2}\beta\frac{\partial\chi}{\partial\beta},\quad b=32\beta^{2}|J|^{2}\chi^{4}{\sf B}.$ (B.4)
If $b\leq\frac{a}{2}$ then we take $p=1$ and in this case we obtain
$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4}(0,x,y,z)|\geq a-b\geq\frac{a}{2}=\chi^{2}\beta\frac{\partial\chi}{\partial\beta}.$ (B.5)
If instead $b>a/2$, then $ap-bp^{2}$ is maximal at $p=\frac{a}{2b}\in(0,1)$, with maximum $\frac{a^{2}}{4b}$. We use this value of $p$ in (B.3), and obtain
$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4}(0,x,y,z)|\geq\frac{a^{2}}{4b}=\frac{1}{32|J|^{2}{\sf B}}\Big{(}\frac{\partial\chi}{\partial\beta}\Big{)}^{2}.$ (B.6)
Then (1.34) follows by choosing the worst alternative in (B.5)–(B.6). ∎
It remains to prove Proposition B.1. Given a vertex set $A\subset\mathbb{T}_{r}$, we define a probability measure $\mathbf{P}^{A}$ on the set $\Omega$ of currents on $\mathbb{T}_{r}$ by
$\mathbf{P}^{A}[\mathbf{n}]=\mathds{1}_{\partial\mathbf{n}=A}\frac{w(\mathbf{n})}{Z^{A}}\qquad(\mathbf{n}\in\Omega),$ (B.7)
where $Z^{A}=\sum_{\partial\mathbf{n}=A}w(\mathbf{n})$ is the normalisation constant. Given two vertex sets $A_{1},A_{2}$, we define
$\mathbf{P}^{A_{1},A_{2}}=\mathbf{P}^{A_{1}}\otimes\mathbf{P}^{A_{2}},$ (B.8)
and we write $\mathbf{E}^{A_{1},A_{2}}$ for expectation with respect to $\mathbf{P}^{A_{1},A_{2}}$.
We write $\mathcal{P}_{2}=\mathcal{P}_{2}({\mathbb{T}_{r}})$ for the set of all unordered pairs of vertices in $\mathbb{T}_{r}$. Let $\mathbf{C}(u)$ denote the vertex cluster of $u\in\mathbb{T}_{r}$ in the percolation configuration constructed from a double current $\mathbf{n}_{1}+\mathbf{n}_{2}$, and let $\mathbf{C}^{b}(u)$ denote the set of pairs $\{s,t\}\in\mathcal{P}_{2}$ such that either $s$ or $t$ belongs to $\mathbf{C}(0)$. Given $\theta\in[0,1]$ and $E\subset\mathcal{P}_{2}$, we introduce the (random) interaction $(J^{E}_{u,v}(\theta))_{\{u,v\}\in\mathcal{P}_{2}}$ defined by
$J^{E}_{u,v}(\theta)=\begin{cases}\theta J_{u,v}&(\{u,v\}\in\mathbf{C}^{b}(0)\cap E)\\ J_{u,v}&(\{u,v\}\not\in\mathbf{C}^{b}(0)\cap E),\end{cases}$ (B.9)
and we write $\langle\cdot\rangle_{\theta}$ for the Ising expectation with the random interaction $J^{E}(\theta)$. Given a subset $G\subset\mathcal{P}_{2}$, we regard $G$ as a graph whose edges consist of the elements of $G$ and whose vertices are the vertices contained in the elements of $G$, and we write $\langle\cdot\rangle^{G}$ for the Ising expectation on the graph $G$ with interaction $J$. By definition, $\langle\cdot\rangle_{1}=\langle\cdot\rangle$ and $\langle\cdot\rangle_{0}=\langle\cdot\rangle^{(\mathbf{C}^{b}(0)\cap E)^{c}}$. According to our convention, the graph $(\mathbf{C}^{b}(0)\cap E)^{c}$ consists of the edges not in $\mathbf{C}^{b}(0)\cap E$, together with the vertices comprising those edges.
We start with a lower bound on $U_{4}$ expressed in terms of $T_{1},T_{2}$ defined for $x,y,z\in\mathbb{T}_{r}$ and $E\subset\mathcal{P}_{2}$ by
$T_{1}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle,$ (B.10)
$T_{2}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.11)
Lemma B.2.
For all $x,y,z\in\mathbb{T}_{r}$, and for any set $E\subset\mathcal{P}_{2}$,
$\frac{1}{2}|{U_{4}}(0,x,y,z)|\geq T_{1}(x,y,z,E)-T_{2}(x,y,z,E).$ (B.12)
Proof.
By [2, Proposition 5.1], for all $x,y,z\in\mathbb{T}_{r}$,
$\frac{1}{2}|U_{4}(0,x,y,z)|$ $=\langle\sigma_{0}\sigma_{x}\rangle\langle\sigma_{y}\sigma_{z}\rangle\mathbf{P}^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)\neq\varnothing]$
$=\langle\sigma_{0}\sigma_{x}\rangle\langle\sigma_{y}\sigma_{z}\rangle\Big{(}1-\mathbf{P}^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]\Big{)}.$ (B.13)
We modify the definition (3.5) by defining, for $H\subset\mathcal{P}_{2}$ and for $A,B$ subsets of vertices in the graph $H$,
$Z^{A,B}_{H}[\;\cdot\;]=\sum_{\begin{subarray}{c}\partial\mathbf{n}_{1}=A\\ \partial\mathbf{n}_{2}=B\end{subarray}}w(\mathbf{n}_{1})w(\mathbf{n}_{2})\mathds{1}\{(\mathbf{n}_{1},\mathbf{n}_{2})\in\cdot\},$ (B.14)
where the sums over $\mathbf{n}_{1},\mathbf{n}_{2}$ are over currents on $H$ and the dot represents any set of pairs of currents. When $H$ is all of $\mathcal{P}_{2}$, we omit the subscript $H$. Then (B) can be written as
$\frac{1}{2}|U_{4}(0,x,y,z)|$ $=\langle\sigma_{0}\sigma_{x}\rangle\langle\sigma_{y}\sigma_{z}\rangle-\frac{Z^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]}{Z^{\varnothing,\varnothing}}.$ (B.15)
To prepare for an application of the switching lemma in the proof of Lemma B.3, we wish to arrange that one of the two currents be sourceless (as they are in $T_{1}$ and $T_{2}$).
This can be achieved by the classical manoeuvre of peeling off the cluster of 00, by conditioning on $\mathbf{C}^{b}(0)$:
$Z^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]$ $=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}_{H}[\mathbf{C}^{b}(0)=H]Z^{\varnothing,yz}_{H^{c}}$
$=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}[\mathbf{C}^{b}(0)=H]\langle\sigma_{y}\sigma_{z}\rangle^{H^{c}}=Z^{0x,\varnothing}\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle^{\mathbf{C}^{b}(0)^{c}}].$ (B.16)
For every $E\subset\mathcal{P}_{2}$, it follows from the second Griffiths inequality that
$\langle\sigma_{y}\sigma_{z}\rangle^{\mathbf{C}^{b}(0)^{c}}\leq\langle\sigma_{y}\sigma_{z}\rangle^{{(\mathbf{C}^{b}(0)\cap E)^{c}}}=\langle\sigma_{y}\sigma_{z}\rangle_{0}.$ (B.17)
Therefore,
$Z^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]$ $\leq Z^{0x,\varnothing}\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle_{0}]=Z^{\varnothing,\varnothing}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle_{0}].$ (B.18)
We insert this into (B.15) and obtain
$\frac{1}{2}|U_{4}(0,x,y,z)|\geq\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle_{1}-\langle\sigma_{y}\sigma_{z}\rangle_{0}].$ (B.19)
By the Fundamental Theorem of Calculus,
$\langle\sigma_{y}\sigma_{z}\rangle_{1}-\langle\sigma_{y}\sigma_{z}\rangle_{0}$ $=\int_{0}^{1}\partial_{\theta_{1}}\langle\sigma_{y}\sigma_{z}\rangle_{\theta_{1}}\mathrm{d}\theta_{1}$
=$=\sum_{\{u,v\}\in E}\beta J_{u,v}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta_{1}}\mathrm{d}\theta_{1}.$ (B.20)
The desired inequality (B.12) then follows from
$\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta_{1}}=\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle-\int^{1}_{\theta_{1}}\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta$ (B.21)
and Fubini’s theorem. ∎
In combination with Lemma B.2, the bounds on $T_{1}$ and $T_{2}$ obtained in the next two lemmas prove Proposition B.1. These bounds use the dilution trick, which consists of picking $E=\omega$ with $\omega$ distributed according to an independent Bernoulli$(p)$ bond percolation on $\mathcal{P}_{2}$, and averaging over $E$. We denote the law and expectation of $\omega$ by $\mathbb{P}_{p}$ and $\mathbb{E}_{p}$, respectively. The dilution method is comparable to the second moment method (see e.g. [4, Lemma 4.4]) but it is more linear in the sense that it allows for resummation. This facilitates working with averaged quantities.
Lemma B.3.
For all $\beta>0$ and for all $p\in[0,1]$,
$\sum_{x,y,z\in\mathbb{T}_{r}}\mathbb{E}_{p}[T_{1}(x,y,z,\omega)]\geq p\chi^{2}\beta\frac{\partial\chi}{\partial\beta}.$ (B.22)
Proof.
By definition,
$T_{1}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle$ (B.23)
$=\frac{\beta}{2}\sum_{u,v\in{\mathbb{T}_{r}}}\mathds{1}[\{u,v\}\in E]J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle.$ (B.24)
We first apply the switching lemma (Lemma 3.2) to see that
$\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\geq\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[0\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\>}u]=\langle\sigma_{0}\sigma_{u}\rangle\langle\sigma_{u}\sigma_{x}\rangle.$ (B.25)
This enables us to take the percolation expectation and obtain
$\mathbb{E}_{p}[T_{1}(x,y,z,\omega)]\geq\frac{\beta p}{2}\sum_{u,v\in\mathbb{T}_{r}}J_{u,v}\langle\sigma_{0}\sigma_{u}\rangle\langle\sigma_{u}\sigma_{x}\rangle\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle.$ (B.26)
Summation over the torus then gives
$\sum_{x,y,z\in\mathbb{T}_{r}}\mathbb{E}_{p}[T_{1}(x,y,z,\omega)]\geq\frac{\beta p}{2}\sum_{y,z,u,v\in{\mathbb{T}_{r}}}J_{u,v}\langle\sigma_{0}\sigma_{u}\rangle\chi\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle.$ (B.27)
By the change of variables $z=z^{\prime}+y$, $u=u^{\prime}+y$, and $v=v^{\prime}+y$, the right-hand side becomes
$\frac{\beta p}{2}\chi\sum_{y,z^{\prime},u^{\prime},v^{\prime}\in{\mathbb{T}_{r}}}J_{u^{\prime},v^{\prime}}\langle\sigma_{0}\sigma_{u^{\prime}+y}\rangle\langle\sigma_{0}\sigma_{z^{\prime}};\sigma_{u^{\prime}}\sigma_{v^{\prime}}\rangle=\beta p\chi^{2}\frac{\partial\chi}{\partial\beta},$ (B.28)
and the proof is complete. ∎
Lemma B.4.
For all $\beta>0$ and for all $p\in[0,1]$,
$\sum_{x,y,z\in\mathbb{T}_{r}}\mathbb{E}_{p}[T_{2}(x,y,z,\omega)]\leq 16p^{2}\beta^{2}|J|^{2}\chi^{4}{\sf B}.$ (B.29)
Proof.
By definition,
$T_{2}(x,y,z,E)=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.30)
An explicit computation gives
$\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}=\beta\sum_{\{s,t\}\in E}J_{s,t}\mathds{1}[s\textup{ or }t\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta},$ (B.31)
where
$\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta}=\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\sigma_{s}\sigma_{t}\rangle_{\theta}-\langle\sigma_{u}\sigma_{v}\rangle_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{s}\sigma_{t}\rangle_{\theta}-\langle\sigma_{s}\sigma_{t}\rangle_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}.$ (B.32)
When the edges $uv$ and $st$ are the same edge, the first term on the right-hand side of (B.32) vanishes, as $\sigma_{u}\sigma_{v}\sigma_{s}\sigma_{t}=1$. In this case, we have $\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{u}\sigma_{v}\rangle_{\theta}\leq 0.$ Therefore,
$T_{2}(x,y,z,E)$ $\leq\frac{\beta^{2}}{4}\sum_{\begin{subarray}{c}u,v,s,t\in{\mathbb{T}_{r}}\\ \{u,v\}\neq\{s,t\}\end{subarray}}\mathds{1}[\{u,v\},\{s,t\}\in E]J_{u,v}J_{s,t}\langle\sigma_{0}\sigma_{x}\rangle$
×$\qquad\times\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\text{or }v\in\mathbf{C}(0)]\mathds{1}[s\text{ or }t\in\mathbf{C}(0)]\int_{0}^{1}\theta\,\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta}\,\mathrm{d}\theta\bigg{]}.$ (B.33)
It is convenient to relabel $u=u_{1}$, $v=u_{2}$, $t=t_{1}$, $s=t_{2}$. For the integral over $\theta$, we use a minor modification [^1] In [6, Proposition 5.2], the Ising model is studied with an external magnetic field $h$. For the random current representation, this involves the introduction of a ghost vertex $g$ with spin $\sigma_{g}=1$, and [6, Proposition 5.2] gives a version of (B.34) in which the vertex $z$ is equal to the ghost. The proof of [6, Proposition 5.2] adapts to our setting with minor modifications. of [6, Proposition 5.2], followed by the monotonicity of the two-point function in $\theta$, to get
$\langle\sigma_{y}\sigma_{z};\sigma_{u_{1}}\sigma_{u_{2}};\sigma_{t_{1}}\sigma_{t_{2}}\rangle_{\theta}\leq\sum_{i,j=1}^{2}2\langle\sigma_{y}\sigma_{u_{i}}\rangle\langle\sigma_{u_{\rho(i)}}\sigma_{t_{j}}\rangle\langle\sigma_{t_{\rho(j)}}\sigma_{z}\rangle+(y\Leftrightarrow z),$ (B.34)
where $\rho(1)=2$, $\rho(2)=1$, and $(y\Leftrightarrow z)$ denotes the sum with the roles of $y$ and $z$ interchanged. This allows us to bound the integral using $\int_{0}^{1}\theta\mathrm{d}\theta=1/2$. For the remaining expectation in (B), we use [4, Proposition A.3], which leads to
$\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u_{1}\textup{ or}u_{2}\in\mathbf{C}(0),t_{1}\textup{ or }t_{2}\in\mathbf{C}(0)]\leq\sum_{i,j=1}^{2}\langle\sigma_{0}\sigma_{u_{i}}\rangle\langle\sigma_{u_{i}}\sigma_{t_{j}}\rangle\langle\sigma_{t_{j}}\sigma_{x}\rangle+(0\Leftrightarrow x),$ (B.35)
where $(0\Leftrightarrow x)$ denotes the sum with the roles of 00 and $x$ interchanged.
We insert the above two inequalities into (B), average over $E=\omega$, and then sum over $x,y,z\in{\mathbb{T}_{r}}$. The sums over $x,y,z$ produce $\chi^{3}$. Then, by translation invariance, we can shift the origin to one of $u_{1},u_{2},t_{1},t_{2}$, and then sum over the former origin to get another factor of $\chi$. The remaining sums are of three kinds and can all be written as convolutions. In terms of the torus convolution defined in (2.9), this produces
$\sum_{x,y,z\in\mathbb{T}_{r}}\mathbb{E}_{p}[T_{2}(x,y,z,\omega)]$
$\qquad\leq\frac{\beta^{2}p^{2}}{4}(\frac{1}{2})(2)(4)\chi^{4}\Big{(}4\lVert J\star\tau^{2}\star J\rVert_{1}+8\lVert J\rVert_{1}(\tau\star J\star\tau)(0)+4(J\star\tau\star J\star\tau)(0)\Big{)},$ (B.36)
where the factor 2 is from (B.34) and the first factor 4 is from $(0\Leftrightarrow x)$ and $(y\Leftrightarrow z)$. We then apply Young’s convolution inequality to see that
$\lVert J\star\tau^{2}\star J\rVert_{1}$ $\leq\lVert J\rVert_{1}^{2}\lVert\tau^{2}\rVert_{1},$ (B.37)
$\lVert J\rVert_{1}(\tau\star J\star\tau)(0)$ $\leq\lVert J\rVert_{1}^{2}\lVert\tau\rVert_{2}^{2},$ (B.38)
$(J\star\tau\star J\star\tau)(0)$ $\leq\lVert J\rVert_{1}^{2}\lVert\tau\rVert_{2}^{2}.$ (B.39)
These right-hand sides are all equal to $|J|^{2}{\sf B}(\beta)$. Altogether, this gives an upper bound $16\beta^{2}p^{2}\chi^{4}\lvert J\rvert^{2}{\sf B}$ for the right-hand side of (B), which concludes the proof. ∎
Appendix C Percolation
The conjectures for the Ising model in Section 1.3 have counterparts for Bernoulli bond (or site) percolation on $\mathbb{Z}^{d}$ in dimensions $d\geq 6$, which we discuss now.
Let $W^{*}$ denote the Brownian excursion of length $1$, let $\Psi(x)=\mathbb{E}\exp[x\int_{0}^{1}W^{*}(t)\mathrm{d}t]$ denote the moment generating function of the Brownian excursion area, let
$F(x,s)=\frac{1}{6}x^{3}-\frac{s}{2}x^{2}+\frac{s^{2}}{2}x,$ (C.1)
and define the conjectured percolation profile
$f_{\rm perc}(s)=\int_{0}^{\infty}x^{2}\mathrm{d}\sigma_{s},\qquad\mathrm{d}\sigma_{s}=\frac{1}{\sqrt{2\pi}}x^{-5/2}\Psi(x^{3/2})e^{-F(x,s)}\mathrm{d}x.$ (C.2)
On the torus, we write $\tau^{\mathbb{T}_{r}}_{p}(0,x)=\mathbb{P}^{\mathbb{T}_{r}}_{p}(0\xleftrightarrow{}x)$ for the percolation two-point function, and $\chi^{\mathbb{T}_{r}}(p)=\sum_{x\in\mathbb{T}_{r}}\tau^{\mathbb{T}_{r}}_{p}(0,x)$ for the susceptibility. On the infinite lattice $\mathbb{Z}^{d}$, we denote the two-point function by $\tau_{p}$ and the critical value by $p_{c}$.
Conjecture C.1 (Universal profile for percolation).
For $d>6$, with window scale $w_{r}=a_{d}r^{-d/3}$ for suitably chosen $a_{d}>0$, and for $s\in\mathbb{R}$, as $r\to\infty$,
$\chi^{\mathbb{T}_{r}}(p_{c}+sw_{r})$ $\sim{\rm const}_{d}\,f_{\rm perc}(s)r^{d/3},$ (C.3)
$\tau^{\mathbb{T}_{r}}_{p_{c}+sw_{r}}(0,x)-\tau_{p_{c}}(0,x)$ $\sim{\rm const}_{d}\,f_{\rm perc}(s)r^{-2d/3}.$ (C.4)
We also conjecture that (C.3) and (C.4) hold for percolation under free boundary conditions (i.e., only using bonds in $\Lambda_{r}$), after $p_{c}$ is replaced by an effective critical value $p_{c,r}=p_{c}+v_{r}$ with $v_{r}\asymp r^{-2}$. If this is true, then since $v_{r}$ is larger than the window $w_{r}$, the two FBC and PBC windows would not overlap.
Upper and lower bounds consistent with (C.3) and (C.4) are proved in [28] for spread-out percolation in dimensions $d>6$, but asymptotic formulas have not been proved and the role of the profile has not been demonstrated. For $d=6$, nothing close to Conjecture C.1 has been proved, though we expect that the profile still applies after appropriate logarithmic corrections to $w_{r}$, $v_{r}$, and to the factors $r^{d/3}$ and $r^{-2d/3}$ in (C.3)–(C.4).
Logarithmic corrections for $d=6$ are discussed in the physics literature, e.g., in [1, 38], on the basis of connections with the $q\to 1$ limit of the $q$-state Potts model and $\varphi^{3}$ field theory. Mathematically, the connection between percolation and $\varphi^{3}$ theory remains mysterious, and to our knowledge there are no rigorous results in this direction. Indeed, the very definition of the $\varphi^{3}$ model is problematic, since $e^{\varphi^{3}}$ is not integrable. Nevertheless, there are indications of a $\varphi^{3}$ structure in the mathematical work on percolation: the percolation diagram in Figure 1 has vertices of degree 3, as do the tree-graph diagrams of [8], and the lace expansion diagrams of [24]. Also, the $x^{3}$ in the function $F$ in $\mathrm{d}\sigma_{s}$ is reminiscent of $\varphi^{3}$.
The conjectured profile arises by analogy with the Erdős–Rényi random graph, where the properly rescaled cluster size (without taking expectation) is known to converge in distribution to a random variable described by the Brownian excursion [9]. The limiting random variable is also characterised by a point process [29]. The measure $\sigma_{s}$ is the intensity of that point process and is taken from [29, Theorem 4.1]. The point process describes cluster sizes, in the sense that
$n^{-2k/3}\sum_{i}\lvert C_{i}\rvert^{k}\Rightarrow\int_{0}^{\infty}x^{k}\mathrm{d}\sigma_{s}\qquad(k\geq 2)$ (C.5)
in distribution [9]. The $k=2$ case gives the properly rescaled susceptibility and identifies $f_{\rm perc}(s)$ in (C.3). To our knowledge, the convergence of the rescaled susceptibility has not been proved, not even for the Erdős–Rényi random graph.
Acknowledgements
We thank Hugo Duminil-Copin, Trishen Gunaratnam, and Alexis Prévost for useful discussions. GS thanks the Section de Mathématiques of the University of Geneva for kind hospitality during a visit when this work was initiated. The work of YL and GS was supported in part by NSERC of Canada. The work of RP was supported by the Swiss National Science Foundation, the NCCR SwissMAP, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 757296).

### v2

@@
 We consider the Ising model on a $d$-dimensional discrete torus of volume $r^{d}$, in dimensions $d>4$ and for large $r$, in the vicinity of the infinite-volume critical point $\beta_{c}$.
-We prove that for $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ (with a suitable constant) the susceptibility is bounded above and below by multiples of $r^{d/2}$, and that the two-point function has a “plateau” in the sense that it decays like $|x|^{-(d-2)}$ when $|x|$ is small relative to the volume but for larger $|x|$ it levels off to a constant value of order $r^{-d/2}$.
+We prove that for $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ (with a suitable constant) the susceptibility is bounded above and below by multiples of $r^{d/2}$.
+Additionally, again for $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$, the two-point function has a “plateau”: it decays like $|x|^{-(d-2)}$ when $|x|$ is small relative to the volume, but for larger $|x|$, it levels off to a constant value of order $r^{-d/2}$.
 We also prove that at $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ the renormalised coupling constant is nonzero, which implies a non-Gaussian limit for the average spin.
@@
 (The claim for $\tau^{\Lambda_{r}}_{\beta_{c}}(0,x)$ is for $x$ not too close to the boundary of the box.) The constant term $r^{-d/2}$ in $\tau^{\mathbb{T}_{r}}_{\beta_{c}}(0,x)$ is referred to as the plateau, and it is responsible for the larger susceptibility for PBC compared to FBC.
-In recent years, the plateau for PBC has been proven to exist for simple random walk in dimensions $d>2$ [40, 15], self-avoiding walk for $d>4$ [40, 30], and percolation for $d>6$ [27, 28].
-A general theory of the effect of FBC vs PBC in the setting of the weakly-coupled hierarchical $n$-component $|\varphi|^{4}$ model in dimensions $d\geq 4$ was developed in [33, 37], and we comment further on this below.
-For the Ising model in dimensions $d>4$, the absence of a plateau for FBC was proven in [13].
+In recent years, the plateau for PBC has been proven to exist for simple random walk in dimensions $d>2$ [43, 16], self-avoiding walk for $d>4$ [43, 30], (spread-out) percolation for $d>6$ [27, 28], and (spread-out) lattice trees and lattice animals for $d>8$ [32].
+A theory of the effect of FBC vs PBC in the setting of the weakly-coupled hierarchical $n$-component $|\varphi|^{4}$ model in dimensions $d\geq 4$ was developed in [35, 40], and we comment further on this below.
+A general theory is outlined in [31].
+For the Ising model in dimensions $d>4$, the absence of a plateau for FBC is an immediate consequence of (1.6) and the monotonicity of $\tau^{\Lambda_{r}}_{\beta_{c}}(0,x)$ in $r$.
+A matching lower bound for FBC is proved in [14].
 For PBC, incomplete steps were taken in [39] towards proving existence of the plateau.
@@
 We emphasise that we do not use the lace expansion in any way, and our results do not require a small parameter.
-The input from the lace expansion used to establish plateau theorems for self-avoiding walk and percolation in [40, 30, 27, 28] is replaced here by the near-critical bounds (1.6)–(1.7).
+The input from the lace expansion used to establish plateau theorems for self-avoiding walk, percolation, and lattice trees and lattice animals in [43, 30, 27, 28, 32] is replaced here by the near-critical bounds (1.6)–(1.7).
 1.2 Main results
@@
 Theorem 1.1.
-Let $d>4$. There are constants $c_{i}>0$, depending on $d$ but not on $\beta$, such that
+Let $d>4$. There are constants $c_{i}>0$, depending on $d$ but not on $\beta$ or $r$, such that
 $\tau_{\beta}^{\mathbb{T}_{r}}(0,x)\leq\tau_{\beta}(0,x)+c_{1}\frac{\chi(\beta)}{r^{d}}$ (1.9)
@@
 The results of [35, 40] are proved for the weakly-coupled $n$-component $|\varphi|^{4}$ model on the $d$-dimensional hierarchical lattice for $d\geq 4$ and for all $n\geq 1$, but, on the basis of universality, the results obtained there for the $1$-component model are conjectured to apply also to all $1$-component spin models in the universality class of the nearest-neighbour Ising model.
-An antecedent of the conjectures goes back to [12] (see also [43, Chapter 32]).
+An antecedent of the conjectures goes back to [13] (see also [46, Chapter 32]).
 The conjectures present predictions for a specific and universal non-Gaussian limit for the average field, and precise asymptotic amplitudes for the susceptibility, the plateau, and the renormalised coupling constant, throughout a critical window of width $r^{-d/2}$ (with a logarithmic correction for $d=4$) centred at $\beta_{c}$.
@@
 $f(s)=\int_{\mathbb{R}}x^{2}\mathrm{d}\sigma_{s},\qquad\mathrm{d}\sigma_{s}=\frac{e^{-\frac{1}{4}x^{4}+\frac{s}{2}x^{2}}\mathrm{d}x}{\int_{\mathbb{R}}e^{-\frac{1}{4}x^{4}+\frac{s}{2}x^{2}}\mathrm{d}x}.$ (1.17)
-The function $f$ is strictly monotone increasing, and satisfies $f(s)\sim|s|^{-1}$ as $s\to-\infty$ and $f(s)\sim s$ as $s\to\infty$ (see [33, Section 1.5.2] where $f_{1}(s)=f(-s)$).
+The function $f$ is strictly monotone increasing, and satisfies $f(s)\sim|s|^{-1}$ as $s\to-\infty$ and $f(s)\sim s$ as $s\to\infty$ (see [35, Section 1.5.2] where $f_{1}(s)=f(-s)$).
 The following conjecture proposes a precise non-Gaussian limit for the average field $S_{r}$, and a precise role for the universal profile $f$ in the critical finite-size scaling of the susceptibility, the two-point function, and the renormalised coupling constant for the Ising model.
 The conjecture is a special case ($n=1$) of conjectures from [35, 40] for $n$-component models for all $n\geq 0$ ($n=0$ corresponds to the self-avoiding walk).
-The profile $f$ also arises in the Curie–Weiss model; see [18, Theorem 3.2].
+The profile $f$ also arises in the Curie–Weiss model; a recent example is [10, Theorem 1.1].
 Conjecture 1.5 (Non-gaussian limit and universal profile).
@@
 $\tau^{\mathbb{T}_{r}}_{\beta_{c}+sw_{r}}(0,x)-\tau_{\beta_{c}}(0,x)\sim h_{r}^{2}f(s)=b_{d}^{2}\,f(s)\times\begin{cases}(\log r)^{1/2}r^{-2}&(d=4)\\ r^{-d/2}&(d>4).\end{cases}$ (1.21)
-As observed in [33, Section 1.5.3], the right-hand side of (1.20) is strictly monotone increasing, and has limits 00 and $2$ as $s\to-\infty$ and $s\to+\infty$ respectively.
+As observed in [35, Section 1.5.3], the right-hand side of (1.20) is strictly monotone increasing, and has limits 00 and $2$ as $s\to-\infty$ and $s\to+\infty$ respectively.
 Thus, assuming (1.20), the renormalised coupling constant varies monotonically over its possible range $[0,2]$ as the window is traversed.
@@
 $\beta_{c,r}=\beta_{c}+v_{r},\qquad v_{r}\asymp\begin{cases}(\log r)^{1/3}r^{-2}&(d=4)\\ r^{-2}&(d>4),\end{cases}$ (1.23)
-such that the statements of Conjecture 1.5 holds as stated when $\beta_{c}$ is replaced by $\beta_{c,r}$.
+such that the statements of Conjecture 1.5 hold as stated when $\beta_{c}$ is replaced by $\beta_{c,r}$.
 In particular, we conjecture that the plateau phenomenon also occurs for free boundary conditions at the effective critical point.
@@
 Proposition 1.7.
-Let $d\geq 1$, $r\geq 3$, and $x\in{\mathbb{T}_{r}}$. Then $\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq T_{\beta}^{(r)}(x).$
+Let $d\geq 1$, $r\geq 3$, $\beta>0$, and $x\in{\mathbb{T}_{r}}$. Then $\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq T_{\beta}^{(r)}(x).$
 Proposition 1.7 suggests decomposing $\tau_{\beta}^{\mathbb{T}_{r}}(x)$ as
@@
 Proof of Corollary 1.2.
-The upper bound follows from (1.9), $\tau_{\beta}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ by (1.6), and the fact that $\chi(\beta_{*})\asymp r^{-d/2}$ from (1.5).
+The upper bound follows from (1.9), $\tau_{\beta}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ by (1.6), and the fact that $\chi(\beta_{*})\asymp r^{d/2}$ from (1.5).
 For the lower bound, when $\lVert x\rVert_{\infty}\geq c_{5}$, we first observe that the lower bound (1.7) applies at $\beta=\beta_{*}$ provided that $r/2\leq c_{0}(\beta_{c}-\beta^{*})^{-1/2}$, i.e.
@@
 Proposition 1.11.
-Let $d\geq 1$. For all $\beta>0$ and for all $r\geq 3$,
-$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4,\beta}^{\mathbb{T}_{r}}(0,x,y,z)|\geq\frac{1}{128d^{2}}\cdot\min\Big{\{}\beta(\chi^{\mathbb{T}_{r}})^{2}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta},\frac{1}{{\sf B}}\Big{(}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta}\Big{)}^{2}\Big{\}}.$ (1.34)
+Let $d\geq 1$. For all $r\geq 3$ and for all $\beta>0$,
+$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4,\beta}^{\mathbb{T}_{r}}(0,x,y,z)|\geq\frac{1}{128d^{2}}\cdot\min\Bigl{\{}\beta(\chi^{\mathbb{T}_{r}})^{2}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta},\frac{1}{{\sf B}}\Big{(}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta}\Big{)}^{2}\Big{\}}.$ (1.34)
 Proof of Theorem 1.4.
@@
 Theorem 2.1.
-Let $d\geq 1$. For any $x\in{\mathbb{T}_{r}}$,
+Let $d\geq 1$, $r\geq 3$, and $\beta>0$. For any $x\in{\mathbb{T}_{r}}$,
 $\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq T_{\beta}^{(r)}(x),$ (2.7)
@@
 Corollary 2.2.
-Let $d\geq 1$. For any $x\in{\mathbb{T}_{r}}$,
-$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq 2\Big{(}T_{\beta}^{(r)}\star\big{(}T_{\beta}^{(r)}\big{)}^{2}[T_{\beta}^{(r)}-(\tau_{\beta}\circ\pi)]\star T_{\beta}^{(r)}\Big{)}(x).$ (2.10)
+Let $d\geq 1$, $r\geq 3$, and $\beta>0$. For any $x\in{\mathbb{T}_{r}}$,
+$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq 2\Big{(}T_{\beta}^{(r)}\star\big{[}\big{(}T_{\beta}^{(r)}\big{)}^{2}[T_{\beta}^{(r)}-(\tau_{\beta}\circ\pi)]\big{]}\star T_{\beta}^{(r)}\Big{)}(x).$ (2.10)
 Proof.
@@
 $\sum_{y}\sum_{z}\tau_{\beta}(z)\tau_{\beta}(y-z)\big{(}T_{\beta}^{(r)}-\tau_{\beta}\big{)}(y-z)T_{\beta}^{(r)}(y-z)T_{\beta}^{(r)}(x-y).$ (2.11)
-Since every point in ${\mathbb{Z}^{d}}$ can be written uniquely in the form $\Lambda_{r}+r{\mathbb{Z}^{d}}$, we can write
+Since every point in ${\mathbb{Z}^{d}}$ can be written uniquely as a point in $\Lambda_{r}+r{\mathbb{Z}^{d}}$, we can write
 $x=\check{x},\quad y=\check{y}+ru,\quad z=\check{z}+rv\qquad(\check{x},\check{y},\check{z}\in\Lambda_{r},\ u,v\in{\mathbb{Z}^{d}}),$ (2.12)
-so the two sums become four sums over $\check{y},\check{z},u,v$.
+so the two sums over $y,z$ become four sums over $\check{y},\check{z},u,v$.
 Since $T_{\beta}^{(r)}$ is periodic, we have $T_{\beta}^{(r)}(y-z)=T_{\beta}^{(r)}(\check{y}-\check{z})$ and $T_{\beta}^{(r)}(x-y)=T_{\beta}^{(r)}(\check{x}-\check{y})$, independent of $u$.
@@
 The counterpart of $T_{\beta}^{(r)}-\tau_{\beta}^{\mathbb{T}_{r}}$ is denoted by $\psi-\psi^{\mathbb{T}}$ in [43, 28, 30].
-For weakly or strictly self-avoiding walk, the bounds [40, (6.36)] and [30, (4.20)] can be written as
+For weakly or strictly self-avoiding walk, the bounds [43, (6.36)] and [30, (4.20)] can be written as
 $T\star[\delta\cdot(T-\tau\circ\pi)]\star T,$ (2.14)
@@
 Also, for $\beta\leq\beta_{c}-c_{4}r^{-d/2}$, by (1.5) we have $r^{-d/2}\chi(\beta)\lesssim c_{4}^{-1}$, and we can make this as small as desired by taking $c_{4}$ large enough.
-With (2.18), this shows that the multiplier of ${\chi(\beta)}/r^{d}$ in (2.17) can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ and $c_{4},c_{5}$ large.
+With (2.18), this shows that the factor multiplying ${\chi(\beta)}/r^{d}$ in (2.17) can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ and $c_{4},c_{5}$ large.
 In particular, it can be made smaller than $\frac{1}{2}c_{T}$, which completes the proof.
@@
 In this section, we introduce a new coupling between Ising models on ${\mathbb{T}_{r}}$ and on large finite subsets of $\mathbb{Z}^{d}$.
-The coupling uses the random current representation (see [16] for an introduction) and a percolation exploration process.
+The coupling uses the random current representation (see [17] or [38, Chapter 2] for an introduction) and a percolation exploration process.
 To derive useful bounds, we develop a new switching lemma between currents on $\mathbb{T}_{r}$ and on large finite subsets of $\mathbb{Z}^{d}$, in Lemma 3.4.
@@
 where
-$w_{\beta}(\mathbf{n})=\prod_{xy\in E}\dfrac{\beta^{\mathbf{n}_{xy}}}{\mathbf{n}_{xy}!},\qquad Z^{A}_{\mathcal{G},\beta}=\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n}).$ (3.3)
+$w_{\beta}(\mathbf{n})=\prod_{e\in E}\dfrac{\beta^{\mathbf{n}_{e}}}{\mathbf{n}_{e}!},\qquad Z^{A}_{\mathcal{G},\beta}=\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n}).$ (3.3)
 For a set of currents $\mathcal{E}\subset\Omega_{\mathcal{G}}$, we also define
 $Z^{A}_{\mathcal{G},\beta}[\mathcal{E}]=\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n})\mathds{1}\{\mathbf{n}\in\mathcal{E}\}.$ (3.4)
+Observe that, by definition, $Z_{\mathcal{G},\beta}^{A}[\Omega_{\mathcal{G}}]=Z^{A}_{\mathcal{G},\beta}$.
 A current $\mathbf{n}$ induces a bond percolation configuration on $\mathcal{G}$ defined by $(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in E}$.
@@
 For a current $\mathbf{n}$ on $\mathcal{G}$ and $x,y\in V$, we write $x\xleftrightarrow{\mathbf{n}\>}y$ if there exists a sequence of points $x_{0}=x,x_{1},\ldots,x_{k}=y$ such that $x_{i}x_{i+1}\in E$ and $\mathbf{n}_{x_{i}x_{i+1}}>0$ for all $0\leq i<k$.
-We will use the following minor extension of the switching lemma of [21, 2] (see also [16, Lemma 7]).
 Given two graphs $\mathcal{G},\mathcal{H}$, two vertex sets $A\subset\mathcal{G}$, $B\subset\mathcal{H}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}\times\Omega_{\mathcal{H}}$, we write
@@
 For two sets $S_{1},S_{2}$, we write $S_{1}\Delta S_{2}=(S_{1}\setminus S_{2})\cup(S_{2}\setminus S_{1})$ for their symmetric difference.
+For two currents $\mathbf{n}$ and $\mathbf{m}$, we write $\mathbf{n}\leq\mathbf{m}$ when $\mathbf{n}_{e}\leq\mathbf{m}_{e}$ for each edge $e$, and we write $\binom{\mathbf{m}}{\mathbf{n}}=\prod_{e\in E}\binom{\mathbf{m}_{e}}{\mathbf{n}_{e}}$ for the product of the binomial coefficients.
+We will use the following minor extension of the switching lemma of [21, 2] (see also [17, Lemma 7]).
 Lemma 3.2 (Switching lemma).
@@
 For a current $\mathbf{n}$ on $\mathcal{G}$, let $\mathbf{n}_{|\mathcal{H}}$ denote its restriction to edges of $\mathcal{H}$.
-Then, for any $A\subset V$, $B\subset V^{\prime}$, $x,y\in V^{\prime}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}$,
-$Z^{A,B}_{\mathcal{G},\mathcal{H}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}$ $=Z^{A\Delta\{x,y\},B\Delta\{x,y\}}_{\mathcal{G},\mathcal{H}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}.$ (3.6)
+Then, for any $A,B\subset V$, $x,y\in V^{\prime}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}$,
+$Z^{A,B}_{\mathcal{G},\mathcal{G}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}$ $=Z^{A\Delta\{x,y\},B\Delta\{x,y\}}_{\mathcal{G},\mathcal{G}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}.$ (3.6)
+It is an immediate consequence of (3.6) that
+$Z^{A,B}_{\mathcal{G},\mathcal{G}}\Big{[}x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}\leq Z^{A\Delta\{x,y\}}_{\mathcal{G}}Z^{B\Delta\{x,y\}}_{\mathcal{G}}.$ (3.7)
 Proof.
@@
 By the change of variables $\mathbf{m}=\mathbf{n}_{1}+\mathbf{n}_{2}$ and $\mathbf{n}=\mathbf{n}_{2}$, the left-hand side of (3.6) can be written as
-$\sum_{\partial\mathbf{m}=A\Delta B}w(\mathbf{m})\mathds{1}\{\mathbf{m}\in\mathcal{E}\}\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.7)
-We view $\mathbf{m},\mathbf{n}$ as multigraphs $\mathcal{M},\mathcal{N}$, and write $\partial\mathcal{N}\subset V^{\prime}$ for the vertices with an odd number of incident edges in $\mathcal{N}$.
+$\sum_{\partial\mathbf{m}=A\Delta B}w(\mathbf{m})\mathds{1}\{\mathbf{m}\in\mathcal{E}\}\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{G}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.8)
+We view $\mathbf{m},\mathbf{n}$ as multigraphs $\mathcal{M},\mathcal{N}$, and write $\partial\mathcal{N}\subset V$ for the vertices with an odd number of incident edges in $\mathcal{N}$.
 From this viewpoint, the sum over $\mathbf{n}$ above exactly counts the number of subgraphs $\mathcal{N}$ of $\mathcal{M}$ with $\partial\mathcal{N}=B$.
@@
 Thus,
-$\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}=\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\Delta\{x,y\}\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.8)
+$\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{G}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}=\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{G}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\Delta\{x,y\}\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.9)
 Undoing the change of variables produces the right-hand side of (3.6). ∎
@@
   • Edges have two possible colourings: white (explored and closed) or black (explored and open). We let $\mathfrak{W}$, $\mathfrak{B}$ be the corresponding evolving sets.
-The construction of $\mathcal{C}(0)$ is as follows. We fix an ordering of the vertices in $\mathcal{G}$.
+The construction of $\mathcal{C}(0)$ is as follows. We fix an (arbitrary) ordering of the vertices in $\mathcal{G}$.
   1. Step 0
@@
 On the other hand, if $(B,W)$ is a valid output of the algorithm and $\omega$ is an element of the right-hand side, then running the exploration on $\omega$ will produce $\mathfrak{B}=B,\>\mathfrak{W}=W$, so the $\supset$ inclusion follows.
-3.3 A coupled cluster exploration
+3.3 A coupled cluster exploration: Proof of (2.7)
 We now apply the exploration introduced in Section 3.2 to a well-chosen percolation configuration induced by random currents.
 This allows us to perform switching between a current on a large finite subgraph $\mathcal{G}\subset\mathbb{Z}^{d}$ and a current on $\mathbb{T}_{r}$.
-Let $\mathbf{m}$ be a current on $\mathcal{G}$ and $\mathbf{n}$ be a current on ${\mathbb{T}_{r}}$. We write $\hat{\mathbf{n}}$ for the lift of $\mathbf{n}$ to $\mathcal{G}$, defined by
-$\hat{\mathbf{n}}(e)=\mathbf{n}(\pi e)\qquad(e\in\mathcal{G}),$ (3.10)
-where $\pi:{\mathbb{Z}^{d}}\to{\mathbb{T}_{r}}$ is the canonical projection map.
+To proceed, we first extend the projection $\pi:\mathbb{Z}^{d}\to\mathbb{T}^{d}$ defined previously, and introduce the notion of a lift from the torus to $\mathbb{Z}^{d}$.
+Projections.
+Given a finite or infinite subgraph $\mathcal{G}=(V,E)$ of $\mathbb{Z}^{d}$, we define its projection $\pi^{\prime}\mathcal{G}=(\pi^{\prime}V,\pi^{\prime}E)$ to be the graph on $\mathbb{T}^{d}$ with vertex set $\pi^{\prime}V=\{\pi x:x\in V\}$ and edge set $\pi^{\prime}E=\{\pi^{\prime}e:e\in E\}$, where for an edge $e=\{x,y\}\in E$ we set $\pi^{\prime}e=\{\pi x,\pi y\}$.
+Since the projection $\pi:\mathbb{Z}^{d}\to\mathbb{T}^{d}$ agrees with the projection $\pi^{\prime}$ applied to a graph consisting of a single vertex, we drop the prime and write simply $\pi$ for the graph projection.
+Thus $\pi$ can act on vertices, on edges, and on graphs.
+When $S$ is a subgraph of ${\mathbb{Z}^{d}}$ whose vertex set does not contain any equivalent vertices, the restriction $\pi_{|S}$ of $\pi$ to $S$ gives an isomorphism between the graphs $S$ and $\pi S$.
+We denote its inverse by $\pi_{|S}^{-1}$.
+Lifts.
+Any function $F$ defined on the vertices and/or edges of the torus $\mathbb{T}^{d}$ can be lifted to the function $\hat{F}$ defined on the vertices/edges of $\mathbb{Z}^{d}$ by $\hat{F}=F\circ\pi$.
+In particular, given a current $\mathbf{n}$ on $\mathbb{T}^{d}$ and a subgraph $\mathcal{G}$ of $\mathbb{Z}^{d}$, the lift $\hat{\mathbf{n}}$ of $\mathbf{n}$ is the current on $\mathcal{G}$ defined by
+$\hat{\mathbf{n}}=\mathbf{n}\circ\pi,\quad\text{i.e.,}\quad\hat{\mathbf{n}}(e)=\mathbf{n}(\pi e)\qquad(e\in\mathcal{G}).$ (3.11)
+Let $\mathbf{m}$ be a current on $\mathcal{G}$ and $\mathbf{n}$ be a current on ${\mathbb{T}_{r}}$.
 The current $\mathbf{m}+\hat{\mathbf{n}}$ induces a percolation configuration $(\mathds{1}_{(\mathbf{m}+\hat{\mathbf{n}})_{e}>0})_{e\in E}$ on $\mathcal{G}$.
@@
 Lemma 3.3 (Properties of the exploration).
-Let $d\geq 1$, $r\geq 3$, $x\in\mathbb{T}_{r}$, and fix finite $\mathcal{G}\supset\Lambda_{2r^{d}}$.
+Let $d\geq 1$, $r\geq 3$, $x\in\mathbb{T}_{r}$, and fix a finite graph $\mathcal{G}\supset\Lambda_{2r^{d}}$.
 Let $\mathbf{m}\in\Omega_{\mathcal{G}}$, and let $\mathbf{n}\in\Omega_{\mathbb{T}_{r}}$ with $\partial\mathbf{n}=\{0,x\}$.
@@
   2. (ii)
-    $\hat{\mathbf{n}}_{e}=0$ for every representative of $e\in\mathfrak{W}$.
+    If $e\in\mathfrak{W}$ then $\hat{\mathbf{n}}_{e^{\prime}}=0$ for every edge $e^{\prime}$ in $\mathcal{G}$ with $\pi{e^{\prime}}=\pi e$.
   3. (iii)
@@
 ∎
-We use the exploration to switch sources between currents on $\mathcal{G}$ and ${\mathbb{T}_{r}}$.
+We use the exploration to switch sources between currents on $\mathcal{G}$ and ${\mathbb{T}_{r}}$, via the following lemma.
 Lemma 3.4.
@@
 $Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[z\in\mathcal{C}(0)]=Z^{C\Delta\{0,z\},A\Delta\{0,\pi z\}}_{\mathcal{G},\mathbb{T}_{r}}[z\in\mathcal{C}(0)].$ (3.12)
-In particular, for $x\in{\mathbb{T}_{r}}$, $A=\{0,x\}$, and $z=x^{\prime}$ with $\pi z=x$,
+In particular, for $x\in{\mathbb{T}_{r}}$, $C=\varnothing$, $A=\{0,x\}$, and $z=x^{\prime}$ with $\pi z=x$,
 $Z^{\varnothing,0x}_{\mathcal{G},\mathbb{T}_{r}}[x^{\prime}\in\mathcal{C}(0)]=Z^{0x^{\prime},\varnothing}_{\mathcal{G},\mathbb{T}_{r}}[x^{\prime}\in\mathcal{C}(0)].$ (3.13)
 Proof.
+Let $\mathbf{m}$ denote the current on $\mathcal{G}$ and $\mathbf{n}$ denote the current on $\mathbb{T}_{r}$.
+Recall that $\mathcal{C}(0)$ consists of black vertices and edges.
 By conditioning on $(\mathfrak{B},\mathfrak{W})$, we have
 $Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[z\in\mathcal{C}(0)]=\sum_{B\ni z,W}Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[(\mathfrak{B},\mathfrak{W})=(B,W)],$ (3.14)
-where $B\ni z$ means that $z$ is incident to some edge in $B$. Given an admissible pair $(B,W)$ (i.e., realisable by the exploration and such that $B\ni z$), we make the decomposition
+where $B\ni z$ means that $z$ is incident to some edge in $B$.
+Henceforth, we write $\mathcal{E}_{(B,W)}$ for the event $\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$, which is measurable in $(\mathbf{m}+\hat{\mathbf{n}})_{|(B\cup W)}$, as indicated at (3.10).
+We will switch sources in $B\cup W$ (which we view as a graph).
+Given an admissible pair $(B,W)$ (i.e., realisable by the exploration and such that $B\ni z$), we make the decomposition
 $\mathbf{m}$ $=\mathbf{m}_{1}+\mathbf{m}_{2},\qquad$ $\mathbf{m}_{1}(e)=\mathbf{m}(e)\mathds{1}\{e\in B\cup W\},$ (3.15)
-$\mathbf{n}$ $=\mathbf{n}_{1}+\mathbf{n}_{2},\qquad$ $~{}\mathbf{n}_{1}(e)=\mathbf{n}(e)\mathds{1}\{e\in\pi(B)\}.$ (3.15)
+$\mathbf{n}$ $=\mathbf{n}_{1}+\mathbf{n}_{2},\qquad$ $~{}\mathbf{n}_{1}(e)=\mathbf{n}(e)\mathds{1}\{e\in\pi(B\cup W)\}.$ (3.16)
 Then $w(\mathbf{m})=w(\mathbf{m}_{1})w(\mathbf{m}_{2})$ and $w(\mathbf{n})=w(\mathbf{n}_{1})w(\mathbf{n}_{2})$.
-The current $\mathbf{n}_{2}$ is supported on $(\pi(B)\cup\pi(W))^{c}$, because $\mathbf{n}=0$ on $\pi(W)$.
-Also, since no edges connect to grey vertices in the exploration, the projection map $\pi$ restricts to an isomorphism between the graphs $B$ and $\pi(B)$, whose inverse we denote by $\iota:\pi(B)\to B$.
-Let $\tilde{\mathbf{n}}_{1}$ denote the restriction of $\hat{\mathbf{n}}_{1}$ to $B$.
-Then
-$\iota^{-1}(\tilde{\mathbf{n}}_{1})=\mathbf{n}_{1},\quad w(\tilde{\mathbf{n}}_{1})=w(\mathbf{n}_{1}),\quad\iota^{-1}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{1}.$ (3.16)
-Let $\mathcal{E}_{(B,W)}$ denote the event $\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$, which is measurable in $\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1}$ as indicated at (3.9).
+We also define $\tilde{\mathbf{n}}_{1}=\hat{\mathbf{n}}_{|B\cup W}$.
+Since no edges connect to grey vertices in the exploration, the projection map $\pi$ restricts to an isomorphism $\pi_{|B\cup W}$ between the graphs $B\cup W$ and $\pi(B\cup W)$, and
+$\tilde{\mathbf{n}}_{1}=\mathbf{n}_{1}\circ\pi_{|B\cup W},\quad w(\tilde{\mathbf{n}}_{1})=w(\mathbf{n}_{1}),\quad\pi_{|B\cup W}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{1}.$ (3.17)
 By conditioning on $\mathbf{m}_{2}$, $\mathbf{n}_{2}$,
-$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}]=\sum_{\mathbf{m}_{2},\mathbf{n}_{2}}w(\mathbf{m}_{2})w(\mathbf{n}_{2})\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\mathbf{n}_{1}=\partial\mathbf{n}_{2}\Delta A\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\},$ (3.17)
-where we have $0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z$ because $B\ni z$.
-The sum over currents $\mathbf{n}_{1}$ on $\pi(B)$ can be replaced by a sum over currents $\tilde{\mathbf{n}}_{1}$ on $B$, with the source constraint $\partial\tilde{\mathbf{n}}_{1}=\iota(\partial\mathbf{n}_{2}\Delta A)$.
-By Lemma 3.2 applied to the graphs $B\cup W$ and $B$,
-$\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\tilde{\mathbf{n}}_{1}=\iota(\partial\mathbf{n}_{2}\Delta A)\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\}$
-$\hskip 42.67912pt=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\Delta\{0,z\}\\ \partial\tilde{\mathbf{n}}_{1}=[\iota(\partial\mathbf{n}_{2}\Delta A)]\Delta\{0,z\}\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\}.$ (3.18)
-Now, we use $\partial\mathbf{n}_{1}=\iota^{-1}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{2}\Delta A\Delta\{0,\pi z\}$ to convert back to a sum over $\mathbf{n}_{1}$. The result is
-$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}]=Z^{C\Delta\{0,z\},A\Delta\{0,\pi z\}}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}].$ (3.19)
+$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}\cap\{z\in B\}]$
+$\qquad\qquad\qquad=\sum_{\mathbf{m}_{2},\mathbf{n}_{2}}w(\mathbf{m}_{2})w(\mathbf{n}_{2})\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\mathbf{n}_{1}=\partial\mathbf{n}_{2}\Delta A\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\},$ (3.18)
+where the connection $0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z$ is present because $B\ni z$.
+In view of (3.17), we can change the sum over $\mathbf{n}_{1}$ to be a sum over $\tilde{\mathbf{n}}_{1}$.
+Now, since $\tilde{\mathbf{n}}_{1}$ and $\mathbf{m}_{1}$ are both currents on $B\cup W$, we can use Lemma 3.2 (with $\mathcal{H}=B$ and $\mathcal{E}=\mathcal{E}_{(B,W)}$) to switch sources between them.
+After the swiching, we use (3.17) again to change the sum over $\tilde{\mathbf{n}}_{1}$ back to a sum over $\mathbf{n}_{1}$.
+With $C_{1}=\partial\mathbf{m}_{2}\Delta C$ and $A_{1}=\partial\mathbf{n}_{2}\Delta A$, this process is represented by
+$\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\\ \partial\mathbf{n}_{1}=A_{1}\end{subarray}}=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\\ \partial\tilde{\mathbf{n}}_{1}=\pi_{|B\cup W}^{-1}A_{1}\end{subarray}}=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\Delta\{0,z\}\\ \partial\tilde{\mathbf{n}}_{1}=(\pi_{|B\cup W}^{-1}A_{1})\Delta\{0,z\}\end{subarray}}=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\Delta\{0,z\}\\ \partial\mathbf{n}_{1}=A_{1}\Delta\{0,\pi z\}\end{subarray}},$ (3.19)
+where the equalities of summation symbols mean that the corresponding replacements are valid in (3.3). After reading the resulting adaptation of (3.3) from right to left, we obtain
+$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}\cap\{z\in B\}]=Z^{C\Delta\{0,z\},A\Delta\{0,\pi z\}}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}\cap\{z\in B\}].$ (3.20)
 Summation over all admissible pairs $(B,W)$ then yields the desired result. ∎
@@
 $\tau_{\beta}^{\mathbb{T}_{r}}(x)=\frac{Z_{\mathbb{T}_{r}}^{0x}}{Z_{\mathbb{T}_{r}}^{\varnothing}}=\frac{Z_{{\Lambda_{R}},\mathbb{T}_{r}}^{\varnothing,0x}}{Z_{{\Lambda_{R}},\mathbb{T}_{r}}^{\varnothing,\varnothing}}=\frac{Z^{\varnothing,0x}_{{\Lambda_{R}},{\mathbb{T}_{r}}}\Big{[}\bigsqcup_{x^{\prime}\cong x}\{x^{\prime}\in\mathcal{C}(0)\}\Big{]}}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{\varnothing,0x}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[x^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}.$ (3.21)
-Using (3.12) and $\{x^{\prime}\in\mathcal{C}(0)\}\subset\{0\xleftrightarrow{\mathbf{m}}x^{\prime}\}$, we get
-$\tau_{\beta}^{\mathbb{T}_{r}}(x)=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[x^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}\leq\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime}).$ (3.21)
+Using (3.13), and then dropping the event $\{x^{\prime}\in\mathcal{C}(0)\}$, we get
+$\tau_{\beta}^{\mathbb{T}_{r}}(x)=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[x^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}\leq\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z_{\Lambda_{R}}^{0x^{\prime}}Z_{\mathbb{T}_{r}}^{\varnothing}}{Z_{\Lambda_{R}}^{\varnothing}Z_{\mathbb{T}_{r}}^{\varnothing}}=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime}).$ (3.22)
 Then (2.7) follows from the fact that $\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime})\leq\tau_{\beta}(x^{\prime})$ by the second Griffiths inequality. ∎
@@
 The proof proceeds by bounding the terms on the right-hand side of (3.24).
-For each $x^{\prime}$, since $\partial\mathbf{m}=\{0,x^{\prime}\}$, there is an open path in $\mathbf{m}$ from 00 to $x^{\prime}$.
+These terms involve a current $\mathbf{m}$ on $\Lambda_{R}$ with $\partial\mathbf{m}=\{0,x^{\prime}\}$, so there is an open path in $\mathbf{m}$ from 00 to $x^{\prime}$.
 When $x^{\prime}\not\in\mathcal{C}(0)$, every such path must contain a grey site.
-(Note that $x^{\prime}$ might be uncoloured, because the sources do not guarantee that there is an $x^{\prime\prime}\in\mathcal{C}(0)$ with $x^{\prime\prime}\cong x$.) We let $y\in{\Lambda_{R}}$ denote the last grey site on one of the paths, so that $y\xleftrightarrow{\mathbf{m}\>}x^{\prime}$ using only edges in $(\mathfrak{B}\cup\mathfrak{W})^{c}$.
+(Note that $x^{\prime}$ might be uncoloured, because the sources do not guarantee that there is an $x^{\prime\prime}\in\mathcal{C}(0)$ with $x^{\prime\prime}\cong x$.) We let $y\in{\Lambda_{R}}$ denote the last grey site on one of the paths from 00 to $x^{\prime}$, so that $y\xleftrightarrow{\mathbf{m}\>}x^{\prime}$ using only edges in $(\mathfrak{B}\cup\mathfrak{W})^{c}$.
 Also, since $y$ is grey, there exists another representative $y^{\prime}\cong y$, $y^{\prime}\neq y$ such that $y^{\prime}\in\mathcal{C}(0)$.
@@
 $\{x^{\prime}\not\in\mathcal{C}(0)\}\subset\bigcup_{y\in{\Lambda_{R}}}\bigcup_{\begin{subarray}{c}y^{\prime}\cong y\\ y^{\prime}\neq y\end{subarray}}\{y^{\prime}\in\mathcal{C}(0)\}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(\mathfrak{B}\cup\mathfrak{W})^{c}\}.$ (3.26)
-In the following, we regard an edge subset (such as $(\mathfrak{B}\cup\mathfrak{W})^{c}$) as the graph whose edges are those in that subset and whose vertices are the vertices comprising those edges.
+As in the proof of Lemma 3.4, we regard an edge subset (such as $(\mathfrak{B}\cup\mathfrak{W})^{c}$) as the graph whose edges are those in that subset and whose vertices are the vertices comprising those edges.
 Lemma 3.6.
-Let $\mathcal{A}(x^{\prime},y,y^{\prime})=\{y^{\prime}\in\mathcal{C}(0)\}\cap\{y\xleftrightarrow{\mathbf{m}}x^{\prime}\text{ in }(\mathfrak{B}\cup\mathfrak{W})^{c}\}$. Then,
+Let $\mathcal{A}(x^{\prime},y,y^{\prime})=\{y^{\prime}\in\mathcal{C}(0)\}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(\mathfrak{B}\cup\mathfrak{W})^{c}\}$ denote the event on the right-hand side of (3.26).
+Then,
 $Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}(x^{\prime},y,y^{\prime})]\leq Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime}).$ (3.27)
 Proof.
+The left-hand side of the desired inequality involves two currents, but we will need a third for the two-point function on the right-hand side.
 For simplicity, we write $\mathcal{A}=\mathcal{A}(x^{\prime},y,y^{\prime})$.
-Since the last event in (3.25) is a connection outside of $\mathfrak{B}\cup\mathfrak{W}$, we can use it to switch sources with a third independent current without affecting $\mathfrak{B}\cup\mathfrak{W}$, as follows.
-Let $\bm{\ell}$ be an independent current on ${\Lambda_{R}}$ with $\partial\bm{\ell}=\varnothing$.
-We again condition on the event $\mathcal{E}_{(B,W)}=\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$.
-We decompose
-$\mathbf{m}=\mathbf{m}_{1}+\mathbf{m}_{2},\qquad\mathbf{m}_{1}(e)=\mathbf{m}(e)\mathds{1}\{e\in B\cup W\},$ (3.27)
-and similarly for $\bm{\ell}=\bm{\ell}_{1}+\bm{\ell}_{2}$. By conditioning on $\mathbf{m}_{1}$ and $\bm{\ell}_{1}$,
-$Z^{\varnothing}_{\Lambda_{R}}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]=Z^{\varnothing}_{\Lambda_{R}}\sum_{B,W}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}\cap\mathcal{E}_{(B,W)}]$
-$\quad=\sum_{B\ni y^{\prime},W}Z^{\varnothing,0x^{\prime},\varnothing}_{{\Lambda_{R}},{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(B\cup W)^{c}\}]$
-$\quad=\sum_{B\ni y^{\prime},W}\sum_{\begin{subarray}{c}\bm{\ell}_{1},\mathbf{m}_{1}\\ \partial\mathbf{n}=\varnothing\end{subarray}}w(\bm{\ell}_{1})w(\mathbf{m}_{1})w(\mathbf{n})\mathds{1}_{\mathcal{E}_{(B,W)}}\!\!\!\!\!\!\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}.$ (3.28)
-Since $\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\subset\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}$, Lemma 3.2 applied with $\mathcal{G}=\mathcal{H}=(B\cup W)^{c}$ gives
-$\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\leq\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\Delta\{y,x^{\prime}\}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,y\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}.$ (3.29)
-We insert this into (3.4), and use $\mathds{1}\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}\leq 1$, to get
-$Z^{\varnothing}_{\Lambda_{R}}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]\leq\sum_{B\ni y^{\prime},W}Z^{yx^{\prime},0y,\varnothing}_{{\Lambda_{R}},{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}]=Z^{yx^{\prime}}_{\Lambda_{R}}Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)].$ (3.30)
-The claim then follows from division by $Z^{\varnothing}_{\Lambda_{R}}$ and (3.2). ∎
-In the proof of Lemma 3.5, we use the notion of the backbone of a current [6, Section 4] (more on the backbone can be found in [5, 35]).
+Since the connection joining $y$ and $x^{\prime}$ is a connection outside of $\mathfrak{B}\cup\mathfrak{W}$, we can use it to switch sources with a third independent current without affecting $\mathfrak{B}\cup\mathfrak{W}$, as follows.
+We again condition on the event $\mathcal{E}_{(B,W)}=\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$ and make the decomposition
+$\mathbf{m}=\mathbf{m}_{1}+\mathbf{m}_{2},\qquad\mathbf{m}_{1}(e)=\mathbf{m}(e)\mathds{1}\{e\in B\cup W\}.$ (3.28)
+Let $\bm{\ell}$ be an independent current on $(B\cup W)^{c}$ with $\partial\bm{\ell}=\varnothing$. By conditioning on $\mathbf{m}_{1}$, we have
+$Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]=\sum_{B,W}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}\cap\mathcal{E}_{(B,W)}]$
+$\quad=\sum_{B\ni y^{\prime},W}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(B\cup W)^{c}\}]\frac{Z^{\varnothing}_{(B\cup W)^{c}}}{Z^{\varnothing}_{(B\cup W)^{c}}}$
+$\quad=\sum_{B\ni y^{\prime},W}\sum_{\begin{subarray}{c}\mathbf{m}_{1}\\ \partial\mathbf{n}=\varnothing\end{subarray}}w(\mathbf{m}_{1})w(\mathbf{n})\mathds{1}_{\mathcal{E}_{(B,W)}}\!\!\!\!\!\!\sum_{\begin{subarray}{c}\partial\bm{\ell}=\varnothing\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\frac{1}{Z^{\varnothing}_{(B\cup W)^{c}}}.$ (3.29)
+We then switch sources between $\bm{\ell}$ and $\mathbf{m}_{2}$ on $\mathcal{G}=\mathcal{H}=(B\cup W)^{c}$.
+Since $\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\subset\{y\xleftrightarrow{\bm{\ell}+\mathbf{m}_{2}\>}x^{\prime}\}$, from (3.7) we get
+$\sum_{\begin{subarray}{c}\partial\bm{\ell}=\varnothing\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\leq\sum_{\partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,y\}}w(\mathbf{m}_{2})\sum_{\partial\bm{\ell}=\{y,x^{\prime}\}}w(\bm{\ell}).$ (3.30)
+After insertion of (3.30) into (3.4), the sum over $\bm{\ell}$, together with $1/Z^{\varnothing}_{(B\cup W)^{c}}$, creates a two-point function on $(B\cup W)^{c}$ which is bounded by $\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime})$ by a Griffiths inequality.
+The remaining factors reassemble to provide the bound
+$Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]\leq\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime})\sum_{B\ni y^{\prime},W}Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}]=\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime})Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)],$ (3.31)
+as desired. ∎
+In the proof of Lemma 3.5, we use the notion of the backbone of a current [6, Section 4] (more on the backbone can be found in [5, 37, 38]).
 A current $\mathbf{m}$ on a finite graph $\mathcal{G}$ with sources $\partial\mathbf{m}=\{u,v\}$ must have at least one path from $u$ to $v$ in the percolation configuration induced by $\mathbf{m}$.
@@
 Let $\mathbf{m}\in\Omega_{\mathcal{G}}$ with $\partial\mathbf{m}=\{u,v\}$.
-The backbone of $\mathbf{m}$, denoted $\Gamma_{\mathbf{m}}$, is the unique oriented and edge self-avoiding path from $u$ to $v$, supported on edges with $\mathbf{m}$ odd, that is minimal for $\prec$.
+The backbone of $\mathbf{m}$, denoted $\Gamma_{\mathbf{m}}$, is the unique oriented and edge self-avoiding path from $u$ to $v$, supported on edges $e$ with $\mathbf{m}_{e}$ odd, that is minimal for $\prec$.
 The backbone $\Gamma_{\mathbf{m}}=\{x_{i}x_{i+1}:0\leq i<k\}$ is obtained via the following exploration process:
@@
 Let $\overline{\Gamma}_{\mathbf{m}}$ denote the set of explored edges, i.e. $\Gamma_{\mathbf{m}}$ together with all explored even edges.
-By definition, $\overline{\Gamma}_{\mathbf{m}}$ is determined by $\Gamma_{\mathbf{m}}$ and the ordering $\prec$, and similarly $\Gamma_{\mathbf{m}}$ can be determined from $\overline{\Gamma}_{\mathbf{m}}$ by selecting odd edges in $\overline{\Gamma}_{\mathbf{m}}$ according to $\prec$.
+By definition, $\overline{\Gamma}_{\mathbf{m}}$ is determined by $\Gamma_{\mathbf{m}}$ and the ordering $\prec$, and similarly $\Gamma_{\mathbf{m}}$ can be determined from $\overline{\Gamma}_{\mathbf{m}}$ by selecting the odd edges in $\overline{\Gamma}_{\mathbf{m}}$ according to $\prec$.
 Also, the current $\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}$, defined to be the restriction of $\mathbf{m}$ to the complement $\overline{\Gamma}_{\mathbf{m}}^{c}$, is sourceless.
@@
 The additional torus two-point function appearing in the right-hand side of (3.25) is produced by the switching of (3.4).
-Since $\partial\mathbf{m}=\{y,y^{\prime}\}$ now, we can form the backbone $\Gamma_{\mathbf{m}}$ of $\mathbf{m}$ from $y^{\prime}$ to $y$.
+Since the current $\mathbf{m}$ on $\Lambda_{R}$ now has sources $\partial\mathbf{m}=\{y,y^{\prime}\}$, we can explore the backbone $\Gamma_{\mathbf{m}}$ of $\mathbf{m}$ from $y^{\prime}$ to $y$.
 Let $\overline{\Gamma}_{\mathbf{m}}$ be the set of edges explored during the formation of $\Gamma_{\mathbf{m}}$.
@@
 For this edge, we must have $z\in\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$ and $\mathbf{m}_{zv}>0$.
-Since one of $z,v$ must be in $\Gamma_{\mathbf{m}}$, we see that $z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}}y^{\prime}$ also.
+Since one of $z,v$ must be in $\Gamma_{\mathbf{m}}$, we see that $z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}\>}y^{\prime}$ also.
 This proves (3.34).
@@
 $Z^{yy^{\prime}}_{\Lambda_{R}}[z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}\>}y^{\prime}]\leq\frac{Z^{yy^{\prime},\varnothing}_{{\Lambda_{R}},{\Lambda_{R}}}[z\xleftrightarrow{\mathbf{m}+\mathbf{m}^{\prime}}y^{\prime}]}{Z^{\varnothing}_{{\Lambda_{R}}}}=\frac{Z^{yz,zy^{\prime}}_{{\Lambda_{R}},{\Lambda_{R}}}}{Z^{\varnothing}_{{\Lambda_{R}}}}=Z^{yz}_{{\Lambda_{R}}}\tau_{\beta}^{{\Lambda_{R}}}(z,y^{\prime}).$ (3.39)
-Inserting this yields
+Inserting this into (3.38) yields
 $\frac{Z^{yy^{\prime},0\pi y^{\prime}}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}$ $\leq\sum_{\begin{subarray}{c}z\in{\Lambda_{R}}\end{subarray}}\frac{Z^{\pi z\pi y^{\prime}}_{\mathbb{T}_{r}}}{Z^{\varnothing}_{\mathbb{T}_{r}}}\frac{Z^{yz}_{{\Lambda_{R}}}}{Z^{\varnothing}_{\Lambda_{R}}}\tau_{\beta}^{{\Lambda_{R}}}(z,y^{\prime})\tau_{\beta}^{{\Lambda_{R}}}(0,z)$
@@
 We present two elementary convolution lemmas.
-The first is a special case of [34, Lemma 3.4].
-We include the simple proof for completeness, and as an opportunity to correct an omission in the proof of [34, Lemma 3.4] (repeated in [33, Lemma 4.2]) which however does not change the conclusion of those lemmas.
+The first is a special case of [36, Lemma 3.4].
+We include the simple proof for completeness, and as an opportunity to correct an omission in the proof of [36, Lemma 3.4] (repeated in [35, Lemma 4.2]) which however does not change the conclusion of those lemmas.
 Lemma A.1.
@@
 where the last inequality is due to $b\leq a-1$ since $x_{0}\geq N_{0}\geq 1$. This completes the proof. ∎
-The next lemma is a minor extension of [34, Lemma 3.6].
+The next lemma is a minor extension of [36, Lemma 3.6].
 The definition of $T_{\beta}^{(r)}(x)$ is in (1.24), $\star$ denotes convolution on $\mathbb{T}_{r}$ and * denotes convolution on $\mathbb{Z}^{d}$.
@@
 $\mathbf{P}^{A}[\mathbf{n}]=\mathds{1}_{\partial\mathbf{n}=A}\frac{w(\mathbf{n})}{Z^{A}}\qquad(\mathbf{n}\in\Omega),$ (B.7)
-where $Z^{A}=\sum_{\partial\mathbf{n}=A}w(\mathbf{n})$ is the normalisation constant. Given two vertex sets $A_{1},A_{2}$, we define
+where $Z^{A}=\sum_{\partial\mathbf{n}=A}w(\mathbf{n})$ is the normalisation constant. Given two vertex sets $A_{1},A_{2}$, we define the product measure
 $\mathbf{P}^{A_{1},A_{2}}=\mathbf{P}^{A_{1}}\otimes\mathbf{P}^{A_{2}},$ (B.8)
@@
 We write $\mathcal{P}_{2}=\mathcal{P}_{2}({\mathbb{T}_{r}})$ for the set of all unordered pairs of vertices in $\mathbb{T}_{r}$.
-Let $\mathbf{C}(u)$ denote the vertex cluster of $u\in\mathbb{T}_{r}$ in the percolation configuration constructed from a double current $\mathbf{n}_{1}+\mathbf{n}_{2}$, and let $\mathbf{C}^{b}(u)$ denote the set of pairs $\{s,t\}\in\mathcal{P}_{2}$ such that either $s$ or $t$ belongs to $\mathbf{C}(0)$.
+Let $\mathbf{C}(u)$ denote the vertex cluster of $u\in\mathbb{T}_{r}$ in the percolation configuration constructed from a double current $\mathbf{n}_{1}+\mathbf{n}_{2}$, and let $\overline{\mathbf{C}}(u)$ denote the set of pairs $\{s,t\}\in\mathcal{P}_{2}$ such that either $s$ or $t$ belongs to $\mathbf{C}(u)$.
 Given $\theta\in[0,1]$ and $E\subset\mathcal{P}_{2}$, we introduce the (random) interaction $(J^{E}_{u,v}(\theta))_{\{u,v\}\in\mathcal{P}_{2}}$ defined by
-$J^{E}_{u,v}(\theta)=\begin{cases}\theta J_{u,v}&(\{u,v\}\in\mathbf{C}^{b}(0)\cap E)\\ J_{u,v}&(\{u,v\}\not\in\mathbf{C}^{b}(0)\cap E),\end{cases}$ (B.9)
-and we write $\langle\cdot\rangle_{\theta}$ for the Ising expectation with the random interaction $J^{E}(\theta)$.
+$J^{E}_{u,v}(\theta)=\begin{cases}\theta J_{u,v}&(\{u,v\}\in\overline{\mathbf{C}}(0)\cap E)\\ J_{u,v}&(\{u,v\}\not\in\overline{\mathbf{C}}(0)\cap E).\end{cases}$ (B.9)
+We write $\langle\cdot\rangle_{\theta}$ for the Ising expectation with the random interaction $J^{E}(\theta)$.
 Given a subset $G\subset\mathcal{P}_{2}$, we regard $G$ as a graph whose edges consist of the elements of $G$ and whose vertices are the vertices contained in the elements of $G$, and we write $\langle\cdot\rangle^{G}$ for the Ising expectation on the graph $G$ with interaction $J$.
-By definition, $\langle\cdot\rangle_{1}=\langle\cdot\rangle$ and $\langle\cdot\rangle_{0}=\langle\cdot\rangle^{(\mathbf{C}^{b}(0)\cap E)^{c}}$.
-According to our convention, the graph $(\mathbf{C}^{b}(0)\cap E)^{c}$ consists of the edges not in $\mathbf{C}^{b}(0)\cap E$, together with the vertices comprising those edges.
+By definition, $\langle\cdot\rangle_{1}=\langle\cdot\rangle$ and $\langle\cdot\rangle_{0}=\langle\cdot\rangle^{(\overline{\mathbf{C}}(0)\cap E)^{c}}$.
+According to our convention, the graph $(\overline{\mathbf{C}}(0)\cap E)^{c}$ consists of the edges not in $\overline{\mathbf{C}}(0)\cap E$, together with the vertices comprising those edges.
 We start with a lower bound on $U_{4}$ expressed in terms of $T_{1},T_{2}$ defined for $x,y,z\in\mathbb{T}_{r}$ and $E\subset\mathcal{P}_{2}$ by
 $T_{1}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle,$ (B.10)
-$T_{2}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.11)
+$T_{2}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}\{u\textup{ or }v\in\mathbf{C}(0)\}\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.11)
 Lemma B.2.
@@
 To prepare for an application of the switching lemma in the proof of Lemma B.3, we wish to arrange that one of the two currents be sourceless (as they are in $T_{1}$ and $T_{2}$).
-This can be achieved by the classical manoeuvre of peeling off the cluster of 00, by conditioning on $\mathbf{C}^{b}(0)$:
-$Z^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]$ $=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}_{H}[\mathbf{C}^{b}(0)=H]Z^{\varnothing,yz}_{H^{c}}$
-$=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}[\mathbf{C}^{b}(0)=H]\langle\sigma_{y}\sigma_{z}\rangle^{H^{c}}=Z^{0x,\varnothing}\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle^{\mathbf{C}^{b}(0)^{c}}].$ (B.16)
+This can be achieved by the classical manoeuvre of peeling off the cluster of 00, by conditioning on $\overline{\mathbf{C}}(0)$:
+$Z^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]$ $=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}_{H}[\overline{\mathbf{C}}(0)=H]Z^{\varnothing,yz}_{H^{c}}$
+$=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}[\overline{\mathbf{C}}(0)=H]\langle\sigma_{y}\sigma_{z}\rangle^{H^{c}}=Z^{0x,\varnothing}\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle^{\overline{\mathbf{C}}(0)^{c}}].$ (B.16)
 For every $E\subset\mathcal{P}_{2}$, it follows from the second Griffiths inequality that
-$\langle\sigma_{y}\sigma_{z}\rangle^{\mathbf{C}^{b}(0)^{c}}\leq\langle\sigma_{y}\sigma_{z}\rangle^{{(\mathbf{C}^{b}(0)\cap E)^{c}}}=\langle\sigma_{y}\sigma_{z}\rangle_{0}.$ (B.17)
+$\langle\sigma_{y}\sigma_{z}\rangle^{\overline{\mathbf{C}}(0)^{c}}\leq\langle\sigma_{y}\sigma_{z}\rangle^{{(\overline{\mathbf{C}}(0)\cap E)^{c}}}=\langle\sigma_{y}\sigma_{z}\rangle_{0}.$ (B.17)
 Therefore,
@@
 $\langle\sigma_{y}\sigma_{z}\rangle_{1}-\langle\sigma_{y}\sigma_{z}\rangle_{0}$ $=\int_{0}^{1}\partial_{\theta_{1}}\langle\sigma_{y}\sigma_{z}\rangle_{\theta_{1}}\mathrm{d}\theta_{1}$
-=$=\sum_{\{u,v\}\in E}\beta J_{u,v}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta_{1}}\mathrm{d}\theta_{1}.$ (B.20)
+=$=\sum_{\{u,v\}\in E}\beta J_{u,v}\mathds{1}\{u\textup{ or }v\in\mathbf{C}(0)\}\int_{0}^{1}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta_{1}}\mathrm{d}\theta_{1}.$ (B.20)
 The desired inequality (B.12) then follows from
@@
 $T_{1}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle$ (B.23)
-$=\frac{\beta}{2}\sum_{u,v\in{\mathbb{T}_{r}}}\mathds{1}[\{u,v\}\in E]J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle.$ (B.24)
+$=\frac{\beta}{2}\sum_{u,v\in{\mathbb{T}_{r}}}\mathds{1}\{\{u,v\}\in E\}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle.$ (B.24)
 We first apply the switching lemma (Lemma 3.2) to see that
@@
 By definition,
-$T_{2}(x,y,z,E)=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.30)
+$T_{2}(x,y,z,E)=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}\{u\textup{ or }v\in\mathbf{C}(0)\}\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.30)
 An explicit computation gives
-$\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}=\beta\sum_{\{s,t\}\in E}J_{s,t}\mathds{1}[s\textup{ or }t\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta},$ (B.31)
+$\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}=\beta\sum_{\{s,t\}\in E}J_{s,t}\mathds{1}\{s\textup{ or }t\in\mathbf{C}(0)\}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta},$ (B.31)
 where
@@
 In this case, we have $\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{u}\sigma_{v}\rangle_{\theta}\leq 0.$ Therefore,
-$T_{2}(x,y,z,E)$ $\leq\frac{\beta^{2}}{4}\sum_{\begin{subarray}{c}u,v,s,t\in{\mathbb{T}_{r}}\\ \{u,v\}\neq\{s,t\}\end{subarray}}\mathds{1}[\{u,v\},\{s,t\}\in E]J_{u,v}J_{s,t}\langle\sigma_{0}\sigma_{x}\rangle$
-×$\qquad\times\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\text{or }v\in\mathbf{C}(0)]\mathds{1}[s\text{ or }t\in\mathbf{C}(0)]\int_{0}^{1}\theta\,\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta}\,\mathrm{d}\theta\bigg{]}.$ (B.33)
+$T_{2}(x,y,z,E)$ $\leq\frac{\beta^{2}}{4}\sum_{\begin{subarray}{c}u,v,s,t\in{\mathbb{T}_{r}}\\ \{u,v\}\neq\{s,t\}\end{subarray}}\mathds{1}\{\{u,v\},\{s,t\}\in E\}J_{u,v}J_{s,t}\langle\sigma_{0}\sigma_{x}\rangle$
+×$\qquad\times\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}\{u\text{ or }v\in\mathbf{C}(0)\}\mathds{1}\{s\text{ or }t\in\mathbf{C}(0)\}\int_{0}^{1}\theta\,\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta}\,\mathrm{d}\theta\bigg{]}.$ (B.33)
 It is convenient to relabel $u=u_{1}$, $v=u_{2}$, $t=t_{1}$, $s=t_{2}$.
@@
 The remaining sums are of three kinds and can all be written as convolutions.
-In terms of the torus convolution defined in (2.9), this produces
+In terms of the torus convolution $\star$ defined in (2.9), this produces
 $\sum_{x,y,z\in\mathbb{T}_{r}}\mathbb{E}_{p}[T_{2}(x,y,z,\omega)]$
 $\qquad\leq\frac{\beta^{2}p^{2}}{4}(\frac{1}{2})(2)(4)\chi^{4}\Big{(}4\lVert J\star\tau^{2}\star J\rVert_{1}+8\lVert J\rVert_{1}(\tau\star J\star\tau)(0)+4(J\star\tau\star J\star\tau)(0)\Big{)},$ (B.36)
-where the factor 2 is from (B.34) and the first factor 4 is from $(0\Leftrightarrow x)$ and $(y\Leftrightarrow z)$. We then apply Young’s convolution inequality to see that
-$\lVert J\star\tau^{2}\star J\rVert_{1}$ $\leq\lVert J\rVert_{1}^{2}\lVert\tau^{2}\rVert_{1},$ (B.37)
+where the factor 2 is from (B.34) and the first factor 4 is from $(0\Leftrightarrow x)$ and $(y\Leftrightarrow z)$. We then apply the Cauchy–Schwarz inequality to see that
+$\lVert J\star\tau^{2}\star J\rVert_{1}$ $=\lVert J\rVert_{1}^{2}\lVert\tau^{2}\rVert_{1},$ (B.37)
 $\lVert J\rVert_{1}(\tau\star J\star\tau)(0)$ $\leq\lVert J\rVert_{1}^{2}\lVert\tau\rVert_{2}^{2},$ (B.38)
@@
 We thank Hugo Duminil-Copin, Trishen Gunaratnam, and Alexis Prévost for useful discussions.
+We also thank two anonymous referees for useful comments.
 GS thanks the Section de Mathématiques of the University of Geneva for kind hospitality during a visit when this work was initiated.